### v1

# An upper bound of the silhouette validation metric for clustering

Abstract
The silhouette coefficient summarizes, per observation, cohesion versus separation in [1,1][-1,1]; the average silhouette width ($\mathrm{ASW}$) is a common internal measure of clustering quality where higher values indicate more coveted results. However, the dataset-specific maximum of $\mathrm{ASW}$ is typically unknown, and the standard upper limit $1$ is often unattainable. In this work, we derive for each data point in a given dataset a sharp upper bound on its silhouette width. By aggregating these individual bounds, we present a canonical data-dependent upper bound on $\mathrm{ASW}$ that often assumes values well below $1$. The presented bounds can indicate whether individual data points can ever be well placed, enable early stopping of silhouette-based optimization loops, and help answer a key question: How close is my clustering result to the best possible outcome on this specific data? Across synthetic and real datasets, the bounds are provably near-tight in many cases and offer significant enrichment of cluster quality evaluation.
keywords:
Data mining , cluster analysis , internal evaluation metric , silhouette score , average silhouette width , upper bound
[…]
1 Introduction
Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [1]. Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [21]. Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [42].
Among cluster validation indices, the silhouette—also known as the silhouette coefficient, silhouette score, or silhouette index—is one of the most widely used measures for assessing how well each point fits within its assigned cluster relative to the nearest neighboring cluster [74]. The average of this measure across all data points is often referred to as the average silhouette width $\mathrm{ASW}$. $\mathrm{ASW}$ values close to 1 signal well-separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned. Extensive work has examined and refined $\mathrm{ASW}$. [11] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency. More recently, [67] proposed alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
Although $\mathrm{ASW}$ is widely used, its raw values are difficult to interpret across heterogeneous datasets. Intrinsic data characteristics can impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$. Directly comparing $\mathrm{ASW}$ values between such datasets can therefore lead to misguided conclusions.
To address this issue, we derive a data-dependent upper bound on the $\mathrm{ASW}$. This bound supplies essential context: if analysis shows that no partition can exceed, say, an $\mathrm{ASW}$ of 0.30, an empirical result of 0.29 should be regarded as essentially optimal. Such insight prevents wasted effort in pursuit of negligible gains and enables more informed evaluation of clustering quality.
A further obstacle is the combinatorial explosion of the partition space. Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion, including the silhouette index, computationally prohibitive. [11] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets. Our contribution complements their approach by providing a principled and data-aware benchmark, namely a theoretical bound, against which the quality of such locally optimal solutions can be evaluated.
Figure 1: Illustration of the silhouette coefficient calculation for $x_{1}$, showing intra-cluster distances within cluster $C_{1}$ and inter-cluster distances to cluster $C_{2}$.
In general, the major contributions of this paper are as follows:
    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$ to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass. We also derive for each data point an upper bound on its silhouette width. These per-point bounds have the potential to highlight observations that cannot be well-placed in any clustering. The bounds are computed by an $O(n^{2}\!\log n)$ algorithm that is practical for medium– to large-scale datasets. In some scenarios, when combined with empirical results, the proposed $\mathrm{ASW}$ upper bound can serve as proof that a silhouette-optimal solution has actually been found.
    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures and clustering baselines. The bound is tight or near-tight in many cases, diagnoses weak clusterability when no inherent group structure exists, tightens under minimum cluster-size constraints, and enables early stopping in silhouette-based model selection.
    All datasets, preprocessing scripts, bound-computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
The rest of this paper is organized as follows. Section 2 reviews related works. Section 3 introduces preliminaries. Section 4 details the proposed method. Section 5 presents the experimental results. Finally, section 6 summarizes the findings and outlines directions for future research.
2 Related work
Silhouette coefficientEnergy & Power Systems[62][73][93][23][27][91][51][94][40][2][19]Transportation & Mobility[60][86][8][79][37][66][58][56][4]Biomedical & Health[43][78][33][34][9][47][75][44][87][12]Food & Agriculture[25][81][32]Social & Economic[63][72][36][90][5]Materials & Chemistry[69][13][76][6][17]Environment & Earth Sciences[61][38][89][41][10][14][35][39][92][18]Sensor Networks & IoT[49][50][64][65][85]Big Data[70][83][82][16]Cluster Validity & Selection[45][46][11][55][15][30][7]Algorithm Development[48][80][59][71][90][8][79][49][50][22][29][31][28][26][32]Feature & Dim-Reduction[75][24][95][16]
Figure 2: Taxonomy of how the silhouette coefficient is employed across research fields.
The silhouette coefficient [74] is a distance-based internal validity index that quantifies how well each object fits its assigned cluster relative to its nearest alternative. Fig. 1 illustrates the construction. For a point $x_{i}\in C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$ (its average dissimilarity to members of $C_{\ell}$) and nearest-cluster separation by $b(x_{i})$ (its average dissimilarity to the closest rival cluster); see Sec.  3 for formal definitions. These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own.
Aggregating over points gives the average silhouette width (ASW), a scale-free summary of partition quality (Sec.  3). In practice, ASW is often used for internal model selection by choosing the number of clusters $K$ that maximizes it; rule-of-thumb interpretations treat larger ASW as evidence of well-separated, compact clusters, and values near zero or negative as signs of overlap or a poor choice of $K$. Despite numerous extensions—e.g., to categorical data [20], fuzzy clustering [53], hierarchical clustering [63], model-based clustering [35], spectral clustering [87], spherical clustering [18], and optimization-based methods [88]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
3 Preliminaries
Table 1: A toy dataset of five samples with two features.
Point $Att_{1}$ $Att_{2}$
$x_{1}$ 1.0 2.0
$x_{2}$ 2.0 1.0
$x_{3}$ 1.5 2.5
$x_{4}$ 6.0 2.0
$x_{5}$ 6.0 3.0
Figure 3: Scatter plot of the toy dataset.
We consider a dataset of $n$ observations described by $m$ features. Throughout, we fix an integer $n\geq 4$ and index the observations by the set $\{1,\dots,n\}$. For example, Tab. 3 and Fig. 3 present a dataset consisting of five data points, each described by two features. We use this toy dataset for illustrative examples in this section.
When applicable, the samples $\{\mathbf{x}_{i}\}_{i=1}^{n}$ of a dataset can be stored in a matrix of the form
$X\;=\;[x_{ij}]_{1\leq i\leq n,\;1\leq j\leq m}\;\in\;\mathbb{R}^{n\times m},\qquad\mathbf{x}_{i}:=(x_{i1},\dots,x_{im})^{\top}\in\mathbb{R}^{m}.$
We deal with datasets where there is a measure of dissimilarity (or distance) between each pair of data points. All pairwise dissimilarities are contained in the datasets dissimilarity matrix.
Definition 1 (Dissimilarity matrix).
A matrix $\Delta\in\mathbb{R}^{n\times n}$ is a dissimilarity matrix if
  1. (a)
    $\Delta_{ij}\geq 0$ for all $i,j$ ($\Delta$ is non-negative),
  2. (b)
    $\Delta_{ii}=0$ for all $i$ ($\Delta$ has zero diagonal), and
  3. (c)
    $\Delta_{ij}=\Delta_{ji}$ for all $i,j$ ($\Delta$ is symmetric).
The triangle inequality is not required. We additionally exclude matrices containing a row of all zeros.
Figure 4: The first 27 Bell numbers, which enumerate the partitions of an $n$-element set. This rapid growth highlights the combinatorial explosion of clustering possibilities as $n$ increases.
For instance, if the Euclidean distance is used, then $\Delta_{ij}=\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert_{2}$. The dissimilarity matrix corresponding to the dataset shown in Fig. 3 is:
$\Delta_{ij}=\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert_{2},\qquad\Delta=\begin{pmatrix}0&1.414&0.707&5.000&5.099\\ 1.414&0&1.581&4.123&4.472\\ 0.707&1.581&0&4.528&4.528\\ 5.000&4.123&4.528&0&1.000\\ 5.099&4.472&4.528&1.000&0\end{pmatrix}$
We remark that the upper bound proposed in Sec. 4 depends only on the dissimilarity matrix of the dataset; no observation vectors are required. Hence, it is just as data-agnostic as the $\mathrm{ASW}$ itself.
For our purposes, it will be convenient to work with the off-diagonal entries of $\Delta$, sorted row-wise in ascending order. We denote this transformed dissimilarity matrix by $\hat{\Delta}$ ($\hat{\Delta}\in\mathbb{R}^{n\times n-1}$). In other words, we have for every $i$ $\{\Delta_{ij}\}_{j=1}^{n}\setminus\{\Delta_{ii}\}=\{\hat{\Delta}_{ij}\}_{j=1}^{n-1}$ and $\hat{\Delta}_{i1}\leq\hat{\Delta}_{i2}\leq\cdots\leq\hat{\Delta}_{i(n-1)}$. Constructing $\hat{\Delta}$ requires sorting each of the $n$ rows of $\Delta$, which costs $\mathcal{O}(n^{2}\log n)$ time in total.
The sorted off-diagonal dissimilarity matrix for the toy dataset is shown in Tab. 2.
Table 2: Row-wise sorted off-diagonal distances $\hat{\Delta}$ corresponding to the toy dataset.
Point $\hat{\Delta}_{i1}$ $\hat{\Delta}_{i2}$ $\hat{\Delta}_{i3}$ $\hat{\Delta}_{i4}$
$x_{1}$ 0.707 1.414 5.000 5.099
$x_{2}$ 1.414 1.581 4.123 4.472
$x_{3}$ 0.707 1.581 4.528 4.528
$x_{4}$ 1.000 4.123 4.528 5.000
$x_{5}$ 1.000 4.472 4.528 5.099
Now, let us formally define the concept of a clustering of the dataset $\{1,\ldots,n\}$.
Definition 2 (Clustering).
Let $K\in\{2,\ldots,n\}$. A $\mathbf{K}$-clustering of $\{i\}_{i=1}^{n}$ is a set $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ such that $C_{I}\subsetneq\{i\}_{i=1}^{n}$ for any $I$, and
[…]
The total number of clusterings of an $n$-element set grows according to the Bell numbers (Fig. 4), illustrating the combinatorial complexity of the clustering problem.
A popular clustering method is the K-means algorithm, which we will use in Sec. 5. Its objective is to minimize the within-cluster sum of squares (or variance).
Definition 3 ($K$-means [57]).
When the desired number of clusters $K$ is specified, the classical $\mathbf{K}$-means problem seeks
$\operatorname*{argmin}_{\mathcal{C}_{K}}\sum_{I=1}^{K}\;\sum_{i\in C_{I}}\bigl{\lVert}\mathbf{x}_{i}-\boldsymbol{\mu}_{I}\bigr{\rVert}_{2}^{2},\qquad\boldsymbol{\mu}_{I}\coloneqq\frac{1}{|C_{I}|}\sum_{i\in C_{I}}\mathbf{x}_{i}.$
Figure 5: Clustering results using $K$-means on the dataset in Fig. 3.
For instance, Fig. 5 illustrates the result of applying $K$-means clustering to the toy dataset with $K=2$ clusters.
Alternative objectives such as linkage-based agglomerative clustering may be formulated directly in terms of $\Delta$, but the underlying aim remains to form internally cohesive and externally separated groups. For instance, Fig.6 shows a dendrogram obtained from performing hierarchical clustering on the dissimilarity matrix corresponding to the toy dataset.
Figure 6: Hierarchical clustering on the distance matrix of data points in Fig. 3.
We will now define the silhouette, which was introduced in 1987 by P.J. Rousseeuw [74]. We employ similar notation.
Definition 4 (Silhouette).
Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{1,\dots,n\}$ with dissimilarity matrix $\Delta$. The silhouette width for data point $i\in C_{I}$ is 0 if $|C_{I}|=1$; otherwise it is
$s(i|\mathcal{C}_{K},\Delta)\coloneqq\frac{b(i)-a(i)}{\max\{a(i),b(i)\}},$
where
$a(i)\coloneqq\frac{1}{|C_{I}|-1}\sum_{j\in C_{I}\setminus\{i\}}\Delta_{ij},\;\;\;\text{and}\;\;\;b(i)\coloneqq\min_{J\neq I}\left(\frac{1}{|C_{J}|}\sum_{j\in C_{J}}\Delta_{ij}\right).$
The average silhouette width (ASW) is defined as
$ASW(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{n}\sum_{i=1}^{n}s(i|\mathcal{C}_{K},\Delta).$
Note that the silhouette width of $i$ may be written as
$s(i|\mathcal{C}_{K},\Delta)=\begin{cases}1-a(i)/b(i)\text{ if }a(i)<b(i),\\ 0\text{ if }a(i)=b(i),\\ b(i)/a(i)-1\text{ if }a(i)>b(i).\end{cases}$
Consequently, $1\leq s(i\,|\,\mathcal{C}_{K},\Delta)\leq 1$.
To illustrate with an example, the silhouette value of each data point from the clustering results shown in Fig. 5 is displayed in Tab. 3 and Fig. 7. The corresponding $\operatorname{ASW}$ is $0.7512$, indicating well-separated and internally cohesive clusters.
Table 3: Silhouette components for each point in the toy dataset after applying K-means clustering with $K=2$.
$i$ $a(i)$ $b(i)$ $s(i)$
$x_{1}$ 1.061 5.050 0.790
$x_{2}$ 1.498 4.298 0.652
$x_{3}$ 1.144 4.528 0.747
$x_{4}$ 1.000 4.550 0.780
$x_{5}$ 1.000 4.700 0.787
Figure 7: Silhouette plots of clustering results from Fig. 5.
Individual silhouette widths provide a valuable graphical diagnostic (see Fig. 7), but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [54]. A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$. When even that maximum is low, the limitation often stems from the intrinsic structure of the dissimilarity data rather than a suboptimal choice of algorithm. This raises the natural question: given a dissimilarity matrix, what is the highest possible $\mathrm{ASW}$? To discuss this question, let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering.
Definition 5 (Silhouette-optimal clustering).
Let $\Delta$ be the dissimilarity matrix of the dataset $\{1,\ldots,n\}$. We define
$\mathfrak{C}\coloneqq\{\mathcal{C}_{K}:\mathcal{C}_{K}\text{ is a $K$-clustering of $\{1,\ldots,n\}$}\}$
to be the set of all possible clusterings. Any
$\mathcal{C}^{*}\in\operatorname*{argmax}_{\mathcal{C}_{K}\in\mathfrak{C}}\{\mathrm{ASW}(\mathcal{C}_{K},\Delta)\}$
is called a silhouette-optimal clustering.
Once again, recall the toy data (Tab. 3) and its dissimilarity matrix $\Delta$. We evaluate $\operatorname{ASW}(\mathcal{C}_{K},\Delta)$ for every admissible clustering with $K\in\{2,3,4,5\}$ (all set partitions except the trivial K=1K{=}1 case). Among the 51 candidates  [^3] The fifth Bell number $B_{5}=52$ and excluding K=1K{=}1 gives $51$., the maximizer is unique:
$\mathcal{C}^{\ast}\;=\;\bigl{\{}\{x_{1},x_{2},x_{3}\},\;\{x_{4},x_{5}\}\bigr{\}}\quad\text{with}\quad\operatorname{ASW}(\mathcal{C}^{\ast},\Delta)\;=\;0.7512,$
which matches the per-point silhouettes reported in Tab. 3 and Fig. 7. Finally, the best achievable $\operatorname{ASW}$ within each fixed $K$ is depicted in Tab. 4.
Table 4: Optimal $\operatorname{ASW}$ within each $K$ on the toy dataset.
$K$ Best partition Best ASW
2 $\{x_{1},x_{2},x_{3}\}\mid\{x_{4},x_{5}\}$ 0.7512
3 $\{x_{1},x_{3}\}\mid\{x_{2}\}\mid\{x_{4},x_{5}\}$ 0.5173
4 $\{x_{1}\}\mid\{x_{2}\}\mid\{x_{3}\}\mid\{x_{4},x_{5}\}$ 0.3068
5 $\{x_{1}\}\mid\{x_{2}\}\mid\{x_{3}\}\mid\{x_{4}\}\mid\{x_{5}\}$ 0.0000
4 The proposed upper bound of the silhouette
4.1 Overview
We propose a data-dependent ceiling on the ASW that can be computed before any clustering is performed. For each data point, we derive a sharp upper bound of its silhouette width. Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and moreover, there exists a clustering in which the bound is exactly achieved. Averaging these per-point ceilings produces a single, dataset-level upper bound on the ASW. The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and supports optional minimum-cluster-size constraints to make it even sharper.
4.2 Derivation
Consider an arbitrary clustering $\mathcal{C}_{K}$ of the dataset $\{1,\ldots,n\}$ with dissimilarites $\Delta$. We will now formalize the upper bound of the silhouette width for point $i$, i.e. $s(i|\mathcal{C}_{K},\Delta)$.
For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$. The average distance between $i$ and the other data points in the same cluster is not smaller than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$. Similarly, the average distance between $i$ and the $n-|C_{I}|$ points farthest from $i$ is not smaller than the average distance between $i$ and the data points in the neighboring cluster closest to $i$. These two observations are combined to construct a value guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$. We formalize this reasoning via the $\Lambda$-quotient of point $i$.
Definition 6.
Let $\Lambda\in\mathbb{N}$ with $1\leq\Lambda\leq n-1$. The $\Lambda$-quotient of data point $i$ is
$q(i,\Delta,\Lambda)\coloneqq\begin{cases}\dfrac{n-\Lambda}{\Lambda-1}\cdot\dfrac{\sum_{j=1}^{\Lambda-1}\hat{\Delta}_{ij}}{\sum_{j=\Lambda}^{n-1}\hat{\Delta}_{ij}}&\textnormal{if }\Lambda>1,\\[4.30554pt] 1,&\textnormal{if }\Lambda=1.\end{cases}$
Recall the following basic fact.
Lemma 1.
Let $A=\{a_{1},\ldots,a_{N}\}$ be a collection of $N$ real values, and consider a partition of $A$ into subsets $A_{1},\ldots,A_{M}$. Then
$\min_{L}\frac{1}{|A_{L}|}\sum_{a\in A_{L}}a\leq\frac{1}{|A|}\sum_{a\in A}a.$
Proof.
Let $\mu=\frac{1}{|A|}\sum_{a\in A}a$ and $\mu_{L}=\frac{1}{|A_{L}|}\sum_{a\in A_{L}}a$. We have $\mu=\sum_{L=1}^{M}\frac{|A_{L}|}{|A|}\mu_{L}$, and since $\sum|A_{L}|=|A|$, the statement holds. ∎
Proposition 1.
Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{i\}_{i=1}^{n}$ with dissimilarites $\Delta$. Let $i$ represent any data point and let $C_{I}\in\mathcal{C}_{K}$ denote the cluster that contains $i$. Furthermore, let
$f(i,\Delta)\coloneqq\min_{1\leq\Lambda\leq n-1}\{q(i,\Delta,\Lambda)\}$
denote a minimal $\Lambda$-quotient. Then
$s(i|\mathcal{C}_{K},\Delta)\leq 1-f(i,\Delta).$
Proof.
Suppose $q(i,\Delta,2)>1$. Then $(n-2)\hat{\Delta}_{i1}>\sum_{j=2}^{n-1}\hat{\Delta}_{ij}$, which contradicts $\hat{\Delta}_{i1}\leq\hat{\Delta}_{ij}$ for $j\geq 2$. Hence, $0\leq f(i,\Delta)\leq 1$. This proves the statement if $|C_{I}|=1$ and for cases $a(i)\geq b(i)$ when $|C_{I}|>1$. Consider $q(i,\Delta,|C_{I}|)$. It is clear that $\sum_{j=1}^{|C_{I}|-1}\hat{\Delta}_{ij}/(|C_{I}|-1)\leq a(i)$. It remains to argue that $\sum_{j=|C_{I}|}^{n-1}\hat{\Delta}_{ij}/(n-|C_{I}|)\geq b(i)$. By Lemma 1, we have
$\frac{1}{n-|C_{I}|}\sum_{j=|C_{I}|}^{n-1}\hat{\Delta}_{ij}$ $\geq\frac{1}{n-|C_{I}|}\sum_{j\in\bigcup_{J\neq I}C_{J}}\Delta_{ij}$
$\geq\min_{J\neq I}\left(\frac{1}{|C_{J}|}\sum_{j\in C_{J}}\Delta_{ij}\right)=b(i).$
Thus, $f(i,\Delta)\leq q(i,\Delta,|C_{I}|)\leq a(i)/b(i)$ and $1-f(i,\Delta)\geq 1-a(i)/b(i)$. ∎
Taking the mean of all silhouette width upper bounds from Proposition 1 results in the following inequality, which is the main observation of this paper:
$ASW(\mathcal{C}^{*},\Delta)\leq 1-\frac{1}{n}\sum_{i}f(i,\Delta).$ (1)
We remark that the right-hand side in 1 depends only on the underlying data $\Delta$. Crucially, we note that $1-\frac{1}{n}\sum_{i}f(i,\Delta)\leq 1$. Moreover, since any $\Lambda$-quotient naturally corresponds to a $2$-clustering of the dataset, the bound in Proposition 1 is indeed sharp.
Finally, imposing the narrower definition
$f_{\kappa}(i,\Delta)\coloneqq\min_{\kappa\leq\Lambda\leq n-\kappa}\{q(i,\Delta,\Lambda)\}$
for some $\kappa\in\{2,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$. Often, as we shall see, this leads to strictly tighter limits.
4.3 Algorithm
Input: $\Delta\in\mathbb{R}^{n\times n}$;
  // Dissimilarity matrix
1
Input: $\kappa$;
  // Smallest allowed cluster size, $1\leq\kappa\leq\lfloor n/2\rfloor$
2
Output: $A\in\mathbb{R}^{n}$;
  // Array of individual upper bounds
3
4 Initialize array $A\leftarrow[\;]$ ;
5 Compute $\hat{\Delta}$ from $\Delta$;
6 for $i=1,\ldots,n$ do
7 $y\leftarrow\sum_{j=\kappa}^{n-1}\hat{\Delta}_{i,j}$ ;
8 $x\leftarrow 0$ ;
 $q\leftarrow 1$ ;
    // Initialize $\Lambda$-quotient
9 if $\kappa>1$ then
10    $x\leftarrow\sum_{j=1}^{\kappa-1}\hat{\Delta}_{i,j}$ ;
11    $q\leftarrow\frac{x}{\kappa-1}\big{/}\frac{y}{n-\kappa}$;
12
13 for $\Lambda=\kappa+1,\ldots,n-\kappa$ do
14    $x\leftarrow x+\hat{\Delta}_{i,\Lambda-1}$;
15    $y\leftarrow y-\hat{\Delta}_{i,\Lambda-1}$;
16    $q_{\text{candidate}}\leftarrow\dfrac{x}{\Lambda-1}\big{/}\dfrac{y}{n-\Lambda}$;
17    if $q_{\text{candidate}}<q$ then
       $q\leftarrow q_{\text{cand}}$ ;
          // Update minimal $\Lambda$-quotient
18
19
20  Append 1q1-q to AA;
21
return AA
Algorithm 1 Silhouette upper bound (pointwise)
The workflow of the proposed upper bound on the ASW is shown in Fig.8. Following this workflow, Algo. 1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the (restricted) solution space $\min_{I}|C_{I}|\geq\kappa$. The steps are as follows: (1) For each row $i$ of $\Delta$, we sort the off-diagonal entries increasingly to obtain $\hat{\Delta}_{i,1},\hat{\Delta}_{i,2},\ldots$, the nearest-to-farthest distances from $i$. (2) For $i$, we initialize the running sums $x$ and $y$ so that $x/(\kappa-1)$ is its mean distance to its $\kappa-1$ nearest points, and $y/(n-\kappa)$ is its mean distance to its $n-\kappa$ farthest points. (3) Starting from $\Lambda=\kappa+1$, we update these running sums as $\Lambda$ increases, thereby scanning all admissible values $\Lambda\in\{\kappa,\ldots,n-\kappa\}$. For each $\Lambda$, we form the $\Lambda$-quotient $q(i,\Delta,\Lambda)$ and keep the minimum. (4) Finally, we record 1q1-q as the upper bound for $s(i\,|\,\mathcal{C}_{K},\Delta)$. The array $A$ collects these values across all points, and its average equals the dataset-level upper bound in Eq. (1) when $\kappa=1$.
Complexity
Forming $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$. The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall run time is $\mathcal{O}(n^{2}\log n)$. This is slightly slower than computing the $\mathrm{ASW}$; $\mathcal{O}(n^{2})$ [54].
Output summaries
Given $\Delta$ and $\kappa=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $A$ from Algo. 1. Then $\texttt{UB}(\Delta)$ corresponds to the right-hand side of Eq. (1). To reflect the spread of the individual bounds, we denote by $\texttt{minUB}(\Delta)$ and $\texttt{maxUB}(\Delta)$ the minimum and maximum values of $A$, respectively. Choosing $\kappa>1$ imposes the restriction that no cluster is allowed to contain fewer than $\kappa$ objects. Let $\texttt{UB}_{\kappa}(\Delta)$ denote the $\mathrm{ASW}$ upper bound for this constrained solution space.
Remark
The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$; its tightness is data-dependent. Nonetheless, it provides an enlightening piece of information that merits inclusion in analyses.
4.4 Example
Recall the toy data from Tab. 3. We have $n=5$, so $\Lambda\in\{1,2,3,4\}$. The values $q(i,\Delta,\Lambda)$ for all instances are shown in Tab. 5. Consider the K=2K{=}2 clustering from Fig. 5: $C_{1}=\{x_{1},x_{2},x_{3}\}$ and $C_{2}=\{x_{4},x_{5}\}$. Tab. 6 compares each silhouette width with the corresponding upper bound, illustrating that $s(i)\leq 1-f(i,\Delta)$ for all $i$, with equality for $x_{2},x_{4},x_{5}$ and a small strict headroom for $x_{1}$ ($0.026$) and $x_{3}$ ($0.054$). Moreover, we have
$\operatorname{ASW}(\mathcal{C}^{\ast},\Delta)=0.7512\;\leq\;0.7672=1-\frac{1}{n}\sum_{i=1}^{n}f(i,\Delta).$
Table 5: All possible values of $q(i,\Delta,\Lambda)$ for each point $i$ in the toy dataset.
Point $q(\cdot,1)$ $q(\cdot,2)$ $q(\cdot,3)$ $q(\cdot,4)$
$x_{1}$ 1.000 0.184 0.210 0.466
$x_{2}$ 1.000 0.417 0.348 0.531
$x_{3}$ 1.000 0.199 0.253 0.502
$x_{4}$ 1.000 0.220 0.538 0.643
$x_{5}$ 1.000 0.213 0.568 0.654
Table 6: Showing $s(i)\leq 1-f(i,\Delta)$ for each point in the toy dataset.
$i$ $|C_{I}|$ $\Lambda^{\star}=\arg\min_{\Lambda}q(i,\Delta,\Lambda)$ $f(i,\Delta)$ $1-f(i,\Delta)$ $s(i)$
$x_{1}$ 3 2 0.184 0.816 0.790
$x_{2}$ 3 3 0.348 0.652 0.652
$x_{3}$ 3 2 0.199 0.801 0.747
$x_{4}$ 2 2 0.220 0.780 0.780
$x_{5}$ 2 2 0.213 0.787 0.787
Figure 8: Workflow of the proposed upper bound on the $\mathrm{ASW}$.
4.5 Enabling early stopping of optimization loops through the upper bound
We now demonstrate how the data–dependent upper bound introduced in Eq. (1) can be used to accelerate standard ASW based optimization loops. In many practical applications, selecting the number of clusters involves iterating over $K$ and evaluating the output of the clustering algorithm $\mathcal{A}(K)$ in order to maximize the ASW. When the dataset is large, each iteration can be computationally demanding, and exhaustive searches over all candidate $K$ quickly become demanding. By incorporating the upper bound into this procedure, we can enable early stopping: if the current ASW is provably within a user–specified relative tolerance of the global optimum, the loop halts, saving potentially substantial computation while still guaranteeing near–optimal clustering performance within the search space. This type of procedure is described in more detail in Algo. 2.
Input: $\Delta\in\mathbb{R}^{n\times n}$;
  // Dissimilarity matrix
1
Input: $\epsilon$;
  // Relative error tolerance, $0\leq\epsilon\leq 1$
2
Input: $\tau$;
  // Clusterability threshold, $0\leq\tau\leq 1$
3
Input: $\mathcal{A}(K),K_{\max}$;
  // Clustering algorithm, max. no. clusters
4
Output: ASW-optimal clustering of $\Delta$ (within search space)
Compute $\mathrm{UB}(\Delta)$;
  // Apply Algo. 1
5 if $\mathrm{UB}(\Delta)\leq\tau$ then // Abort if dataset is insufficiently clusterable
6 return NULL
7optimal clustering $\leftarrow$ NULL ;
8 optimal ASW $\leftarrow$ -1 ;
9 for $K=2,\ldots,K_{\max}$ do
[…]
11      optimal clustering $\leftarrow\mathcal{A}(K)$ ;
12      optimal ASW $\leftarrow\mathrm{ASW}(\mathcal{A}(K),\Delta)$ ;
13    if $\frac{\mathrm{UB}(\Delta)-\text{optimal ASW}}{\mathrm{UB}(\Delta)}<\epsilon$ then
14       return optimal clustering
15
16
return optimal clustering
Algorithm 2 Early stopping in ASW-based optimization
Workflow
The procedure integrates the proposed upper bound into a standard ASW optimization loop. At each step, the current best clustering is compared to the global upper bound. If the relative gap falls below the user-specified tolerance $\epsilon$, the algorithm terminates early, certifying that no further search can improve ASW by more than $\epsilon\%$ of the true global optimum. In the worst case, the condition is never met, and the loop reverts to the standard exhaustive search over $K=2,\ldots,K_{\max}$, ensuring no loss in clustering quality.
Remark
If minimal cluster size constraints are of interest, the procedure can be modified by replacing the initial call to $\mathrm{UB}(\Delta)$ with the restricted version $\mathrm{UB}_{\kappa}(\Delta)$. This enables early stopping under user-specified requirements. We also emphasise that clusterability in this context is interpreted solely through the lens of the silhouette coefficient. While this provides a useful and practical criterion for our purposes, it is by no means a universal or definitive notion of clusterability—many alternative perspectives exist in the literature.
Recall the toy dataset summarized in Tab. 4. In Sec.  4.4, we established that the corresponding upper bound is $\texttt{UB}(\Delta)=0.7672$. Let $\widehat{S}_{K}$ denote the empirically observed $\mathrm{ASW}$ value for $K$ clusters. For the toy dataset, these values are:
$\widehat{S}_{2}=0.7512,\quad\widehat{S}_{3}=0.5173,\quad\widehat{S}_{4}=0.3068,\quad\widehat{S}_{5}=0.0000.$
The performance metric of interest is the relative error, defined as
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}_{K}}{\mathrm{ASW}(C^{*},\Delta)},$
where $C^{*}$ denotes the clustering achieving the global maximum ASW. Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}_{K}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}_{K}}{\texttt{UB}(\Delta)}.$
We refer to this quantity as the worst-case relative error. Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical $\widehat{S}_{K}$ is guaranteed to be within $\epsilon$ of the true maximum ASW.
Table 7: Early stopping on the toy dataset.
$K$ $\widehat{S}_{K}$ $\texttt{UB}(\Delta)$ worst-case relative error
[…]
As illustrated by Tab. 7, applying Algo. 2 with relative error tolerance $\epsilon=0.05$ for the toy dataset would halt the process already at $K=2$, thereby avoiding unnecessary evaluations for $K=3,4,5$. In this way, the upper bound allows substantial savings while still certifying near-optimal performance.
5 Experiments, results & discussion
5.1 Experimental datasets
Table 8: Datasets used in our experiments and their characteristics. The Metric(s) column indicates the distance metrics used in the experiments.
ID Dataset Name Instances Features Classes Field Metric(s)
1 gene Gene expressions 801 20531 Unknown Genomics Euclidean, correlation
2 ceramic Ceramic samples 88 17 Unknown Archaeology Euclidean
3 relig. texts Religious texts 589 8266 Unknown Humanities Cosine, Euclidean, Jaccard
4 conf. papers Conference papers 5804 11463 Unknown Computer science Cosine, Euclidean, Jaccard
5 wdbc Breast Cancer Wisconsin Diagnostic 569 30 2 Healthcare Euclidean
6 wisc Breast Cancer Wisconsin Original 699 9 2 Healthcare Euclidean
7 ecoli Ecoli 336 7 8 Biology Euclidean
8 wine Wine 178 13 3 Chemistry Euclidean
9 iono Ionosphere 351 34 2 Physics Euclidean
10 zoo Zoo 101 16 7 Biology Euclidean
11 iris Iris 150 4 3 Botany Euclidean
12 segment Image Segmentation 2310 19 7 Computer science Euclidean
13 dermatology Dermatology 366 34 6 Healthcare Euclidean
14 balance-scale Balance Scale 625 4 3 Psychology Euclidean
15 heart-statlog Statlog (Heart) 270 13 2 Healthcare Euclidean
16 thy New Thyroid 215 5 3 Healthcare Euclidean
We evaluate the upper bound by applying it to a variety of datasets, both synthetic and real-life. Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity. (For some parts of the experiments we use the Jaccard distance, which means the data is treated in binary form and similarity is based on the amount of shared attributes.)
An essential part of the evaluation is to establish a tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$. To this end, we adopt an empirical approach: we apply several clustering algorithms to each dataset and take the largest observed $\mathrm{ASW}$ as the lower bound. The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies.
The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
Figure 9: Average silhouette width $\mathrm{ASW}$ vs. number of clusters ($K$). Results on the synthetic dataset 400-64-5-6 for $K$-means, PAM (PAMSIL), and Hierarchical clustering with weighted and single linkage. For each method and each chosen $K$, we compute the resulting $\mathrm{ASW}$. The black dashed line marks the theoretical upper bound $\mathrm{UB}(\Delta)=0.376$; the red dashed line marks the upper bound when a minimum cluster size of $80$ is enforced ($0.282$). The maximum observed $\mathrm{ASW}$ occurs at K=5K=5—a partition into five equal-sized clusters of 80 elements—and is achieved by $K$-means, Hierarchical (weighted linkage), and PAMSIL.
Synthetic datasets
We follow [3] in notation and approach. Synthetic datasets are generated using the make_blobs function from scikit-learn [68]. Each dataset is labeled as n_samplesn_featurescenterscluster_stdn\_samples-n\_features-centers-cluster\_std, indicating the parameter values used to generate the points. For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion.
Real-life datasets
We use sixteen datasets spanning unlabeled (IDs 1–4) and labeled (IDs 5–16) settings (Tab. 8). The unlabeled datasets are named gene expressions, ceramic samples, religious texts, and conference papers. For these, we search over $K$ for different algorithms to maximize $\operatorname{ASW}$. The labeled datasets consist of twelve standard benchmarks across multiple domains-wdbc, wisc, ecoli, wine, iono, zoo, iris, segment, dermatology, balance-scale, heart-statlog, thy-ranging from 150 to 2,310 instances, 4 to 34 features, and 2 to 8 classes. For these datasets, we use the ground-truth class count to set $K$ (e.g., $K=8$ for ecoli), run $K$-medoids to obtain a $K$-clustering, and compare its $\operatorname{ASW}$ with the bound; we also repeat a search over varying $K$ to maximize $\operatorname{ASW}$, mirroring the unlabeled protocol.
Table 9: Experimental results on synthetic data. Each dataset is denoted as n_samples-n_features-centers-cluster_std. For Hierarchical clustering (weighted and single linkage), $K$-means, and $K$-medoids (Hybrid), we report the average silhouette width (ASW); the selected number of clusters $K$ is shown in parentheses. For $K$-medoids, $\dagger$ indicates that PAMSIL was used; datasets with $\geq\!1000$ samples used PAMMEDSIL to reduce runtime. The last three columns summarize the output of Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ upper bound), $\min\mathrm{UB}(\Delta)$, and $\max\mathrm{UB}(\Delta)$ (smallest and largest individual silhouette width upper bound respectively). All results use Euclidean distance.
Dataset Metric
Weighted
$K$-means
$\texttt{UB}(\Delta)$ $\texttt{minUB}(\Delta)$ $\texttt{maxUB}(\Delta)$
400-64-5-6 Euclidean $0.249$ (5) $0.209$ (4) $0.249$ (5) $0.249^{\dagger}$ (5) $0.376$ $0.276$ $0.467$
400-64-2-2 Euclidean $0.673$ (2) $0.673$ (2) $0.673$ (2) $0.673^{\dagger}$ (2) $0.673$ $0.619$ $0.718$
400-128-7-3 Euclidean $0.522$ (7) $0.522$ (7) $0.522$ (7) $0.522^{\dagger}$ (7) $0.566$ $0.514$ $0.619$
1000-161-2-13 Euclidean $0.084$ (2) $0.068$ (2) $0.084$ (2) $0.046$ (3) $0.182$ $0.137$ $0.256$
1000-300-5-2 Euclidean $0.668$ (5) $0.668$ (5) $0.668$ (5) $0.668$ (5) $0.678$ $0.651$ $0.706$
10000-32-20-2 Euclidean $0.626$ (20) $0.626$ (20) $0.626$ (20) $0.626$ (20) $0.774$ $0.681$ $0.847$
10000-1024-20-4 Euclidean $0.417$ (20) $0.417$ (20) $0.417$ (20) $0.417$ (20) $0.454$ $0.425$ $0.488$
Table 10: Experimental results on real unlabeled data. For each dataset (the distance/dissimilarity used is shown in the “Metric” column), we report the average silhouette width (ASW) for Hierarchical clustering (weighted and single linkage), $K$-means, and $K$-medoids (Hybrid); the selected number of clusters $K$ is given in parentheses. For $K$-medoids, $\dagger$ indicates PAMSIL; datasets with $\geq\!1000$ samples used PAMMEDSIL to reduce runtime. When the metric is incompatible with $K$-means (e.g., correlation/cosine or Jaccard), the entry is marked N/A. The last three columns summarize the output of Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ upper bound), $\min\mathrm{UB}(\Delta)$, and $\max\mathrm{UB}(\Delta)$ (smallest and largest individual silhouette width upper bound respectively).
Dataset Metric
Weighted
$K$-means
$\texttt{UB}(\Delta)$ $\texttt{minUB}(\Delta)$ $\texttt{maxUB}(\Delta)$
Gene expressions Correlation $0.345$ (10) $0.332$ (2) N/A $0.395^{\dagger}$ (7) $0.657$ $0.325$ $0.865$
Religious texts Cosine $0.082$ (234) $0.049$ (2) N/A $0.087^{\dagger}$ (23) $0.349$ $0.079$ $0.810$
Conference papers Cosine $0.188$ (2) $0.162$ (2) N/A $0.126$ (2) $0.473$ $0.118$ $1.0$
Religious texts Euclidean $0.846$ (2) $0.846$ (2) $0.428$ (3) $0.846^{\dagger}$ (2) $0.846$ $0.106$ $0.895$
Ceramic samples Euclidean $0.547$ (2) $0.591$ (2) $0.584$ (2) $0.584^{\dagger}$ (2) $0.855$ $0.561$ $0.950$
Conference papers Euclidean $0.418$ (2) $0.418$ (2) $0.067$ (2) $0.384$ (2) $0.439$ $0.227$ $1.0$
Gene expressions Euclidean $0.207$ (2) $0.188$ (2) $0.228$ (6) $0.230^{\dagger}$ (8) $0.419$ $0.174$ $0.624$
Religious texts Jaccard $0.031$ (106) $0.017$ (569) N/A $0.024^{\dagger}$ (24) $0.119$ $0.040$ $0.719$
Conference papers Jaccard $0.160$ (2) $0.160$ (2) N/A $0.160$ (2) $0.170$ $0.038$ $1.0$
Table 11: Early stopping on real data. Early stopping results (Algo. 2) with Euclidean dissimilarity. Best ASW per row is bold; corresponding $K$ is bold.
Dataset
$K$
cand.
$\epsilon$
Best ASW
($K$-means)
Best ASW
($K$-medoids)
UB
Best $K$
($K$-means)
Best $K$
($K$-medoids)
Worst-case
relative error
($K$-means)
Worst-case
relative error
($K$-medoids)
Early stop
($K$-means)
Early stop
($K$-medoids)
gene $[2,15]$ 0.15 0.228 0.230 0.419 6 8 0.46 0.45 false false
ceramic $[2,30]$ 0.35 0.584 0.584 0.855 2 2 0.32 0.32 true true
relig. texts $[2,15]$ 0.15 0.428 0.846 0.846 3 2 0.49 0.0 false true
conf. papers $[2,15]$ 0.15 0.067 0.384 0.439 2 2 0.85 0.13 false true
wdbc $[2,15]$ 0.15 0.345 0.661 0.708 2 2 0.51 0.07 false true
wisc $[2,15]$ 0.15 0.574 0.574 0.844 2 2 0.32 0.32 false false
ecoli $[2,15]$ 0.15 0.398 0.836 0.854 4 2 0.53 0.02 false true
wine $[2,30]$ 0.35 0.285 0.295 0.626 3 5 0.55 0.53 false false
iono $[2,15]$ 0.15 0.293 0.413 0.691 6 2 0.58 0.40 false false
zoo $[2,30]$ 0.35 0.547 0.547 0.839 30 22 0.35 0.35 true true
iris $[2,30]$ 0.35 0.580 0.580 0.878 2 2 0.34 0.34 true true
segment $[2,15]$ 0.15 0.357 0.778 0.929 7 2 0.62 0.16 false false
dermatology $[2,15]$ 0.15 0.277 0.357 0.609 3 2 0.55 0.42 false false
balance-scale $[2,15]$ 0.15 0.219 0.214 0.737 8 8 0.70 0.71 false false
heart-statlog $[2,30]$ 0.35 0.169 0.322 0.593 3 2 0.72 0.46 false false
thy $[2,30]$ 0.35 0.577 0.734 0.835 2 2 0.31 0.12 true true
5.2 Benchmarking algorithms
To cluster the datasets, we employ three popular methods: $K$-means, $K$-medoids, and Hierarchical agglomerative clustering. These methods serve as baselines to contextualize the performance of the proposed upper bound.
Because our goal is to obtain the highest possible $\mathrm{ASW}$, we evaluate multiple values of $K$ and report the one that yields the largest $\mathrm{ASW}$.
$K$-means clustering
We use the implementation of $K$-means from scikit-learn. For each dataset, the algorithm requires the number of clusters $K$.
$K$-medoids clustering
We use the Python package $K$-medoids [77] and, in particular, two medoid-based algorithms: PAMSIL and PAMMEDSIL, both introduced in [52]. PAMSIL aims to directly optimize $\mathrm{ASW}$ but is computationally expensive. PAMMEDSIL is a faster proxy that optimizes a medoid-based version of the silhouette; while expedient, it generally results in a lower $\mathrm{ASW}$ than PAMSIL. To balance accuracy and run time, we use PAMSIL for datasets with $n<1000$ and switch to PAMMEDSIL for $n\geq 1000$. This threshold was chosen empirically to keep computation tractable while preserving $\mathrm{ASW}$ quality.
Hierarchical agglomerative clustering (HAC)
For hierarchical clustering, we use the SciPy implementation [84]. To measure inter-cluster distances when forming the hierarchy, we consider both single and weighted linkage.
Figure 10: Silhouette plots for the 4 real unlabeled datasets. $K$-means clusters; ASW compared against $\texttt{UB}(\Delta)$ and the size-constrained bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size.
Figure 11: Silhouette plots for the 12 real labeled datasets. $K$-medoids with $K$ fixed to the ground truth; ASW compared against $\texttt{UB}(\Delta)$.
Figure 12: Silhouette plots for the 12 real labeled datasets. $K$-medoids with $K$ chosen by unrestricted optimization; ASW compared with $\texttt{UB}(\Delta)$, illustrating the best empirical quality without fixing $K$ to ground truth.
Figure 13: Unlabeled datasets: violin plots of pointwise silhouette upper bounds versus minimum cluster size $\kappa$ (up to $\sim 10\%$ of $n$). The violins summarize distributions across data points (means, i.e. $\mathrm{ASW}$ upper bounds, in red). The upper bounds decline as $\kappa$ increases, showing that size constraints lead to sharper limits.
5.3 Results and discussion
Fig. 9 presents the $\mathrm{ASW}$ obtained by different clustering algorithms on the synthetic dataset 400-64-5-6. K-means achieves its best result with five clusters of equal size ($80$ objects each), reaching an $\mathrm{ASW}$ of $0.249$. Hierarchical clustering with weighted linkage also favors five clusters, whereas single linkage attains its maximum $\mathrm{ASW}$ ($0.209$) at four clusters. K-medoids reaches the same output as K-means and weighted linkage. The theoretical upper bound without restrictions on cluster size ($\kappa=1$) is $0.376$ (black dashed line). When enforcing a minimum cluster size of $\kappa=80$, Algo. 1 yields a reduced upper bound of $0.282$ (red dashed line). Thus, among solutions with no cluster smaller than $80$ elements, the five-cluster partition found by e.g. K-means is guaranteed to be within $12\%$ of the global optimum.
Tab. 5.1 reports the remaining results on the synthetic datasets. For each algorithm, we show the best observed $\mathrm{ASW}$, with the selected number of clusters in parentheses. The last three columns summarize the upper bounds from Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ bound), $\texttt{minUB}(\Delta)$ (smallest individual bound), and $\texttt{maxUB}(\Delta)$ (largest individual bound). For dataset 400-64-2-2, the upper bound matches the achieved $\mathrm{ASW}$ for all algorithms, certifying global optimality. Dataset 1000-161-2-13 has a significantly low upper bound ($\mathrm{UB}(\Delta)=0.182$), implying that no clustering can exceed an $\mathrm{ASW}$ of 0.182; in such cases, alternative quality measures may be more appropriate, or the data may be effectively non-clusterable.
Tab. 5.1 reports results on the real unlabeled datasets and mirrors the layout of Tab. 5.1. By default, we search over $K\in\{2,\ldots,n\}$, where $n$ is the number of samples. When necessary for reasonable run time, however, we limit the search at $K\leq 30$. Entries marked N/A indicate that $K$-means is incompatible with the metric, e.g., correlation, cosine or Jaccard. For Euclidean “Religious texts”, Euclidean “Conference papers”, and Jaccard “Conference papers”, the hierarchical methods are within $0.021$ of the upper bound $\mathrm{UB}(\Delta)$—so any further improvement in ASW over the reported values is at most $0.021$. Several datasets have small $\texttt{minUB}(\Delta)$ (e.g., Cosine “Religious texts” with $\texttt{minUB}(\Delta)=0.079$), certifying that at least one observation cannot achieve a silhouette width exceeding that value under any clustering. In general, when $\texttt{minUB}(\Delta)$ is near zero, it may be reasonable to treat the corresponding observations as noise or remove them from the analysis.
We implement the early-stopping algorithm presented in Sec.  4.5 and examine empirically how it performs on the real datasets. In Tab. 11, we adopt the following convention: For datasets with $n\leq 300$, we explore $K\in\{2,\dots,30\}$ with tolerance $\epsilon=0.35$. For datasets with $n>300$, we restrict the search to $K\in\{2,\dots,15\}$ with tolerance $\epsilon=0.15$. These parameter choices are not canonical, but rather serve to illustrate different practical regimes. The goal is to demonstrate how Algo. 2 behaves under varying $K$ candidates and tolerance, rather than to prescribe a universally optimal setting.
The results indicate that the proposed upper bound can indeed shorten silhouette-based optimization loops in practice under reasonable tolerance choices. In several cases — notably ceramic, religious-texts, conference-papers, wdbc, ecoli, zoo, iris, and thy — the algorithm halted early while still returning an ASW within the specified tolerance. In other cases, the bound was not tight enough compared to the observed output to trigger early stopping, and the procedure effectively reverted to a full search. This mixed outcome highlights both the promise and the limitations of the approach: when the bound interacts favorably with the empirical ASW, it can reduce computation substantially, but there is no guarantee of savings in every scenario.
Fig. 10 visualizes for the real unlabeled datasets the individual upper bounds in the context of the clusterings obtained by K-means (see Tab. 5.1). It also shows how the upper bound is tightened under size constraints. Fig. 11 and Fig. 12 are constructed in the same way, but for the labeled datasets, without size constraints and with K-medoids instead of K-means. The difference between the two of them is that the latter implements a scan of different $K$ values to maximize the empirical ASW whereas the former uses the ground truth labels to stipulate $K$. Fig. 13 also corresponds to the labeled data and illustrates a more rigorous analysis of the distribution of the individual upper bounds in relation to the $\kappa$ parameter of Algo. 1.
We conclude the discussion by highlighting three key roles of the proposed upper bound.
  1. 1.
    Global upper bound. The unrestricted bound $\mathrm{UB}(\Delta)$ is a ceiling on any achievable ASW for the dataset. It can shorten silhouette-based search loops by enabling early stopping; when tight, it certifies that a globally silhouette-optimal clustering has been found. Even when not tight, it provides essential context for interpreting reported ASW values across algorithms and initializations. Moreover, the magnitude of $\mathrm{UB}(\Delta)$ itself can be seen as a proxy for the clusterability of the dataset: a low upper bound suggests that no clustering can yield well-separated and compact groups, whereas a high upper bound indicates the potential for meaningful structure.
  2. 2.
    Restricted upper bound. Analysts often care about a constrained solution space. By enforcing a minimum cluster size via $\kappa$, the sharper bound $\mathrm{UB}_{\kappa}(\Delta)$ reflects application constraints and yields insights that are more relevant.
  3. 3.
    Individual upper bounds. Our proposed pointwise upper bounds quantify the maximum silhouette each observation can attain. Very low values flag observations that cannot be well integrated under any clustering; such points may merit down-weighting, manual inspection, or exclusion as outliers or noise.
6 Conclusion
In this paper, we propose a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a new tool for analyzing clustering quality. While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a much tighter, data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time. This upper bound provides a more meaningful reference value: it often lies well below $1$, reflecting the inherent limitations of the dataset itself.
Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often close-to-sharp and yields valuable insights into the structure and limitations of the data. At the same time, we emphasize that the bound is not guaranteed to be sharp in general; the true global ASW-maximum may fall considerably below the bound. For this reason, the bound should be viewed not as a precise target but as an informative ceiling that frames observed ASW scores in a broader context. Even when loose, the bound can shorten optimization loops, offer a first-order estimate of the datasets clusterability, and highlight points that are inherently difficult to cluster.
From a broader perspective, our main contribution is to establish a proof of concept: that efficiently computable, data-dependent upper bounds on ASW can enrich cluster analysis. We believe this idea opens a promising line of work, where future research should aim at deriving tighter bounds, extending the approach to other clustering validity indices, and exploring additional constraints motivated by application domains. A particularly important direction is to develop theoretical guarantees, for example bounding how far the proposed upper bound can deviate from the true global maximum under certain assumptions on the dataset. This goal could be attained through the derivation of a theoretical and close-to-sharp lower bound on $\mathrm{ASW}(\mathcal{C}^{*},\Delta)$.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement
Hugo Sträng: conceptualization, data curation, methodology, experiment, visualization, writing, editing; Tai Dinh: visualization, writing, editing, validation.
Acknowledgments
The authors thank Hugo Fernström for his careful proofreading and for offering insightful feedback on a previous version. His constant encouragement greatly contributed to the development of this work.

### v2

@@
 Abstract
-The silhouette coefficient summarizes, per observation, cohesion versus separation in [1,1][-1,1]; the average silhouette width ($\mathrm{ASW}$) is a common internal measure of clustering quality where higher values indicate more coveted results.
-However, the dataset-specific maximum of $\mathrm{ASW}$ is typically unknown, and the standard upper limit $1$ is often unattainable.
-In this work, we derive for each data point in a given dataset a sharp upper bound on its silhouette width.
-By aggregating these individual bounds, we present a canonical data-dependent upper bound on $\mathrm{ASW}$ that often assumes values well below $1$.
-The presented bounds can indicate whether individual data points can ever be well placed, enable early stopping of silhouette-based optimization loops, and help answer a key question: How close is my clustering result to the best possible outcome on this specific data?
-Across synthetic and real datasets, the bounds are provably near-tight in many cases and offer significant enrichment of cluster quality evaluation.
+The silhouette coefficient quantifies, for each observation, the balance between within-cluster cohesion and between-cluster separation, taking values in [1,1][-1,1].
+The average silhouette width (ASW) is a widely used internal measure of clustering quality, with higher values indicating more cohesive and well-separated clusters.
+However, the dataset-specific maximum of ASW is typically unknown, and the standard upper limit of 1 is rarely attainable.
+In this work, we derive for each data point a sharp upper bound on its silhouette width and aggregate these to obtain a canonical upper bound on the ASW.
+This bound—often substantially below 1—enhances the interpretability of empirical ASW values by indicating how close a given clustering result is to the best possible outcome on that dataset.
+It can be used to confirm global optimality, guide the evaluation of clustering solutions, and be refined to incorporate minimum cluster-size constraints for greater practical relevance.
+Finally, we extend the framework to establish a corresponding bound for the macro-averaged silhouette.
 keywords:
@@
 1 Introduction
-Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [1].
-Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [21].
-Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [42].
-Among cluster validation indices, the silhouette—also known as the silhouette coefficient, silhouette score, or silhouette index—is one of the most widely used measures for assessing how well each point fits within its assigned cluster relative to the nearest neighboring cluster [74].
+Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [aggarwal2013introduction].
+Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [dinh2025data].
+Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [jain1999data].
+Among the cluster validation indices, the silhouette width (also known as the silhouette coefficient, silhouette score, or silhouette index) is one of the most widely used measures to assess how well each point fits within its assigned cluster relative to its nearest neighboring cluster [ogsilhouette].
 The average of this measure across all data points is often referred to as the average silhouette width $\mathrm{ASW}$.
-$\mathrm{ASW}$ values close to 1 signal well-separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned.
+$\mathrm{ASW}$ values close to 1 signal well‑separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned.
 Extensive work has examined and refined $\mathrm{ASW}$.
-[11] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency.
-More recently, [67] proposed alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
+[batool2021clustering] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency.
+More recently, [john] examined alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
 Although $\mathrm{ASW}$ is widely used, its raw values are difficult to interpret across heterogeneous datasets.
-Intrinsic data characteristics can impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$.
-Directly comparing $\mathrm{ASW}$ values between such datasets can therefore lead to misguided conclusions.
+Intrinsic data characteristics impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$.
+Comparing $\mathrm{ASW}$ values across datasets with different properties can therefore lead to misguided conclusions about cluster quality.
 To address this issue, we derive a data-dependent upper bound on the $\mathrm{ASW}$.
@@
 Such insight prevents wasted effort in pursuit of negligible gains and enables more informed evaluation of clustering quality.
-A further obstacle is the combinatorial explosion of the partition space.
-Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion, including the silhouette index, computationally prohibitive.
-[11] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets.
-Our contribution complements their approach by providing a principled and data-aware benchmark, namely a theoretical bound, against which the quality of such locally optimal solutions can be evaluated.
-Figure 1: Illustration of the silhouette coefficient calculation for $x_{1}$, showing intra-cluster distances within cluster $C_{1}$ and inter-cluster distances to cluster $C_{2}$.
-In general, the major contributions of this paper are as follows:
-    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$ to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass.
-We also derive for each data point an upper bound on its silhouette width.
-These per-point bounds have the potential to highlight observations that cannot be well-placed in any clustering.
-The bounds are computed by an $O(n^{2}\!\log n)$ algorithm that is practical for medium– to large-scale datasets.
-In some scenarios, when combined with empirical results, the proposed $\mathrm{ASW}$ upper bound can serve as proof that a silhouette-optimal solution has actually been found.
-    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures and clustering baselines.
-The bound is tight or near-tight in many cases, diagnoses weak clusterability when no inherent group structure exists, tightens under minimum cluster-size constraints, and enables early stopping in silhouette-based model selection.
-    All datasets, preprocessing scripts, bound-computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
+The biggest obstacle in silhouette optimization is the combinatorial explosion of the solution space as the number of samples increases.
+Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion computationally prohibitive.
+[batool2021clustering] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets.
+Our contribution complements their approach by providing a data-aware benchmark against which the quality of such locally optimal solutions can be evaluated.
+Figure 1: Illustration of the silhouette coefficient calculation.
+The main contributions are as follows:
+    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$, computable in $O(n^{2}\!\log n)$ time, to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass.
+    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures.
+We show that the bound is nearly tight in many of the cases examined.
+    All datasets, preprocessing scripts, bound computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
 The rest of this paper is organized as follows.
@@
 2 Related work
-Silhouette coefficientEnergy & Power Systems[62][73][93][23][27][91][51][94][40][2][19]Transportation & Mobility[60][86][8][79][37][66][58][56][4]Biomedical & Health[43][78][33][34][9][47][75][44][87][12]Food & Agriculture[25][81][32]Social & Economic[63][72][36][90][5]Materials & Chemistry[69][13][76][6][17]Environment & Earth Sciences[61][38][89][41][10][14][35][39][92][18]Sensor Networks & IoT[49][50][64][65][85]Big Data[70][83][82][16]Cluster Validity & Selection[45][46][11][55][15][30][7]Algorithm Development[48][80][59][71][90][8][79][49][50][22][29][31][28][26][32]Feature & Dim-Reduction[75][24][95][16]
+Silhouette coefficientEnergy & Power Systems[ouyang2024k][rodriguez2024novel][zheng2025practical][duarte2021increasing][eskandarnia2022embedded][zhao2023generalized][koutsandreas2024harnessing][zhou2025recognition][hu2023transformer][alrasheed2024building][diaz2022typologies]Transportation & Mobility[mohd2024passengers][wang2021analyzing][bai2023adaptive][sheng2025multi][guo2025adaptive][passarella2024anomaly][ma2025unravelling][li2025clustering][ansari2021spatiotemporal]Biomedical & Health[jamalpour2025high][sharafoddini2021identifying][ghaderi2023self][gill2024symptom][baier2024nocturnal][kim2022pathological][rustam2022data][jha2021apache][xian2025systematic][bhagat2022dpcf]Food & Agriculture[edris2024identifying][singh2025development][gere2023recommendations]Social & Economic[owoo2025hierarchical][quintero2022analysis][gulma2022new][zhang2025two][aonso2024cluster]Materials & Chemistry[pham2025clustering][bi2025comprehensive][sancho2022cluster][bagirov2023novel][consuegra101unsupervised]Environment & Earth Sciences[nourani2022robust][guo2023combining][yang2022zoning][huang2023dominant][balbi2024hierarchical][biabiany2020design][guan2025novel][gupta2025criteria][zhao2023comparison][deng2025on]Sensor Networks & IoT[kotary2020distributed][kotary2021many][panwar2024distributed][panwar2025distributed][waidyanatha2023regularity]Big Data[qaroush2025maximum][thirumoorthy2021hybrid][song2023novel][chander2023dolphin]Cluster Validity & Selection[khan2025standardizing][khan2025standardization][batool2021clustering][lenssen2024medoid][capo2023fast][gagolewski2021are][bagirov2023finding]Algorithm Development[koren2025preprocessing][silva2025automatic][manochandar2020development][qin2025single][zhang2025two][bai2023adaptive][sheng2025multi][kotary2020distributed][kotary2021many][du2024multi][fu2025gsscl][gan2024discrimination][fu2023gc][enriquez2025performance][gere2023recommendations]Feature & Dim‑Reduction[rustam2022data][dwivedi2025novel][zhou2024ocae][chander2023dolphin]
 Figure 2: Taxonomy of how the silhouette coefficient is employed across research fields.
-The silhouette coefficient [74] is a distance-based internal validity index that quantifies how well each object fits its assigned cluster relative to its nearest alternative.
+The silhouette coefficient [ogsilhouette] is a distance-based internal validity index that quantifies how well each object fits into its assigned cluster relative to its nearest alternative.
 Fig.
 1 illustrates the construction.
-For a point $x_{i}\in C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$ (its average dissimilarity to members of $C_{\ell}$) and nearest-cluster separation by $b(x_{i})$ (its average dissimilarity to the closest rival cluster); see Sec.
-3 for formal definitions.
+For a point $x_{i}$ in cluster $C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$—its average dissimilarity to the other members of $C_{\ell}$—and nearest-cluster separation by $b(x_{i})$—its average dissimilarity to the closest neighboring cluster.
 These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own.
-Aggregating over points gives the average silhouette width (ASW), a scale-free summary of partition quality (Sec.
-3).
-In practice, ASW is often used for internal model selection by choosing the number of clusters $K$ that maximizes it; rule-of-thumb interpretations treat larger ASW as evidence of well-separated, compact clusters, and values near zero or negative as signs of overlap or a poor choice of $K$.
-Despite numerous extensions—e.g., to categorical data [20], fuzzy clustering [53], hierarchical clustering [63], model-based clustering [35], spectral clustering [87], spherical clustering [18], and optimization-based methods [88]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
+Aggregating over points gives the ASW, a dataset-level summary of partition quality.
+In practice, ASW is often used for model selection by choosing the number of clusters $K$ that maximizes it.
+Despite numerous extensions—e.g., to categorical data [dinh2019estimating], fuzzy clustering [lasek2024employing], hierarchical clustering [owoo2025hierarchical], model-based clustering [guan2025novel], spectral clustering [xian2025systematic], spherical clustering [deng2025on], and optimization-based methods [yang2024clustering]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
+Other commonly used internal validation indices include the Davies–Bouldin index [dbindex] and the Calinski–Harabasz index [chindex], which also evaluate compactness and separation but rely on alternative formulations of intra- and inter-cluster dispersion.
+In this work, we focus on the silhouette due to its interpretability and widespread adoption across both research and applied domains.
+To account for cluster-size imbalance, we additionally employ the macro-averaged silhouette, defined as the unweighted mean of per-cluster silhouette averages.
+This variant ensures that smaller clusters contribute equally to the overall assessment, providing a more balanced view of clustering quality.
+Silhouette-optimization
+PAMSIL (Partitioning Around Medoids based on SILhouette) is a modification of the classical PAM algorithm [kaufman90], designed to directly optimize the ASW [pamsil].
+Unlike $K$-means or standard $K$-medoids, which minimize within-cluster dissimilarities, PAMSIL explicitly seeks cluster configurations that maximize ASW.
+It achieves this by modifying the SWAP phase of PAM: in each iteration, the algorithm performs the swap that yields the largest improvement in ASW and terminates once no further improvement is possible, ensuring convergence to a local optimum.
+An efficient implementation of PAMSIL is available in the open-source K-Medoids Python package [Schubert2022].
+Moreover, Fatima Batool and Christian Hennig have proposed alternative algorithms that find local ASW optima [batool2021clustering].
+External validation indices
+When ground-truth class labels are available, external validation measures provide a way to quantify how well a clustering agrees with a known reference partition.
+Two of the most commonly used indices are the adjusted Rand index (ARI) [ari] and the adjusted mutual information (AMI) [ami].
+Both adjust their underlying similarity measures for chance agreement, ensuring that random labelings yield expected scores near zero.
+The ARI extends the classical Rand index by comparing all pairs of samples and counting how consistently they are assigned to the same or different clusters in both partitions.
+Its normalization accounts for random overlap, producing values in [1,1][-1,1] with $1$ indicating identical partitions.
+The AMI applies a similar correction to the mutual information (MI) between partitions, normalizing it by the expected MI under random assignments and bounding the result between 0 and $1$.
+Both indices are symmetric, easy to interpret, and widely used in benchmarking studies, where they enable consistent comparison of clustering algorithms across datasets with known class structure.
 3 Preliminaries
-Table 1: A toy dataset of five samples with two features.
-Point $Att_{1}$ $Att_{2}$
-$x_{1}$ 1.0 2.0
-$x_{2}$ 2.0 1.0
-$x_{3}$ 1.5 2.5
-$x_{4}$ 6.0 2.0
-$x_{5}$ 6.0 3.0
-Figure 3: Scatter plot of the toy dataset.
-We consider a dataset of $n$ observations described by $m$ features.
-Throughout, we fix an integer $n\geq 4$ and index the observations by the set $\{1,\dots,n\}$.
-For example, Tab.
-3 and Fig.
-3 present a dataset consisting of five data points, each described by two features.
-We use this toy dataset for illustrative examples in this section.
-When applicable, the samples $\{\mathbf{x}_{i}\}_{i=1}^{n}$ of a dataset can be stored in a matrix of the form
-$X\;=\;[x_{ij}]_{1\leq i\leq n,\;1\leq j\leq m}\;\in\;\mathbb{R}^{n\times m},\qquad\mathbf{x}_{i}:=(x_{i1},\dots,x_{im})^{\top}\in\mathbb{R}^{m}.$
-We deal with datasets where there is a measure of dissimilarity (or distance) between each pair of data points. All pairwise dissimilarities are contained in the datasets dissimilarity matrix.
+Let $n\in\mathbb{N}$ with $n\geq 4$.
+We consider a dataset consisting of $n$ objects indexed by the set $\{1,\ldots,n\}$.
+Each pair of objects is associated with a measure of dissimilarity (or distance), and all pairwise dissimilarities are collected in the dataset’s dissimilarity matrix.
 Definition 1 (Dissimilarity matrix).
@@
   1. (a)
-    $\Delta_{ij}\geq 0$ for all $i,j$ ($\Delta$ is non-negative),
+    $\Delta_{ij}\geq 0$ for all $i,j$ ($\Delta$ is non‑negative),
   2. (b)
@@
     $\Delta_{ij}=\Delta_{ji}$ for all $i,j$ ($\Delta$ is symmetric).
-The triangle inequality is not required. We additionally exclude matrices containing a row of all zeros.
+The triangle inequality is not required. We also exclude matrices containing a row of all zeros.
 Figure 3: The first 27 Bell numbers, which enumerate the partitions of an $n$-element set. This rapid growth highlights the combinatorial explosion of clustering possibilities as $n$ increases.
-For instance, if the Euclidean distance is used, then $\Delta_{ij}=\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert_{2}$. The dissimilarity matrix corresponding to the dataset shown in Fig. 3 is:
-$\Delta_{ij}=\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert_{2},\qquad\Delta=\begin{pmatrix}0&1.414&0.707&5.000&5.099\\ 1.414&0&1.581&4.123&4.472\\ 0.707&1.581&0&4.528&4.528\\ 5.000&4.123&4.528&0&1.000\\ 5.099&4.472&4.528&1.000&0\end{pmatrix}$
-We remark that the upper bound proposed in Sec.
-4 depends only on the dissimilarity matrix of the dataset; no observation vectors are required.
-Hence, it is just as data-agnostic as the $\mathrm{ASW}$ itself.
 For our purposes, it will be convenient to work with the off-diagonal entries of $\Delta$, sorted row-wise in ascending order.
@@
 Constructing $\hat{\Delta}$ requires sorting each of the $n$ rows of $\Delta$, which costs $\mathcal{O}(n^{2}\log n)$ time in total.
-The sorted off-diagonal dissimilarity matrix for the toy dataset is shown in Tab. 2.
-Table 2: Row-wise sorted off-diagonal distances $\hat{\Delta}$ corresponding to the toy dataset.
-Point $\hat{\Delta}_{i1}$ $\hat{\Delta}_{i2}$ $\hat{\Delta}_{i3}$ $\hat{\Delta}_{i4}$
-$x_{1}$ 0.707 1.414 5.000 5.099
-$x_{2}$ 1.414 1.581 4.123 4.472
-$x_{3}$ 0.707 1.581 4.528 4.528
-$x_{4}$ 1.000 4.123 4.528 5.000
-$x_{5}$ 1.000 4.472 4.528 5.099
 Now, let us formally define the concept of a clustering of the dataset $\{1,\ldots,n\}$.
@@
 Let $K\in\{2,\ldots,n\}$. A $\mathbf{K}$-clustering of $\{i\}_{i=1}^{n}$ is a set $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ such that $C_{I}\subsetneq\{i\}_{i=1}^{n}$ for any $I$, and
+  1. (a)
+    $\bigcup_{I=1}^{K}C_{I}=\{1,\dots,n\}$,
+  2. (b)
 […]
+  3. (c)
+    $C_{I}\neq\varnothing\textnormal{ for all }I\in\{1,\ldots,K\}$.
 The total number of clusterings of an $n$-element set grows according to the Bell numbers (Fig. 3), illustrating the combinatorial complexity of the clustering problem.
-A popular clustering method is the K-means algorithm, which we will use in Sec. 5. Its objective is to minimize the within-cluster sum of squares (or variance).
-Definition 3 ($K$-means [57]).
-When the desired number of clusters $K$ is specified, the classical $\mathbf{K}$-means problem seeks
-$\operatorname*{argmin}_{\mathcal{C}_{K}}\sum_{I=1}^{K}\;\sum_{i\in C_{I}}\bigl{\lVert}\mathbf{x}_{i}-\boldsymbol{\mu}_{I}\bigr{\rVert}_{2}^{2},\qquad\boldsymbol{\mu}_{I}\coloneqq\frac{1}{|C_{I}|}\sum_{i\in C_{I}}\mathbf{x}_{i}.$
-Figure 5: Clustering results using $K$-means on the dataset in Fig. 3.
-For instance, Fig. 5 illustrates the result of applying $K$-means clustering to the toy dataset with $K=2$ clusters.
-Alternative objectives such as linkage-based agglomerative clustering may be formulated directly in terms of $\Delta$, but the underlying aim remains to form internally cohesive and externally separated groups.
-For instance, Fig.6 shows a dendrogram obtained from performing hierarchical clustering on the dissimilarity matrix corresponding to the toy dataset.
-Figure 6: Hierarchical clustering on the distance matrix of data points in Fig. 3.
-We will now define the silhouette, which was introduced in 1987 by P.J. Rousseeuw [74]. We employ similar notation.
+We will now define the silhouette width, which was introduced in 1987 by P.J. Rousseeuw [ogsilhouette]. We employ similar notation.
 Definition 3 (Silhouette).
@@
 Consequently, $1\leq s(i\,|\,\mathcal{C}_{K},\Delta)\leq 1$.
-To illustrate with an example, the silhouette value of each data point from the clustering results shown in Fig.
-5 is displayed in Tab.
-3 and Fig.
-7.
-The corresponding $\operatorname{ASW}$ is $0.7512$, indicating well-separated and internally cohesive clusters.
-Table 3: Silhouette components for each point in the toy dataset after applying K-means clustering with $K=2$.
-$i$ $a(i)$ $b(i)$ $s(i)$
-$x_{1}$ 1.061 5.050 0.790
-$x_{2}$ 1.498 4.298 0.652
-$x_{3}$ 1.144 4.528 0.747
-$x_{4}$ 1.000 4.550 0.780
-$x_{5}$ 1.000 4.700 0.787
-Figure 7: Silhouette plots of clustering results from Fig. 5.
-Individual silhouette widths provide a valuable graphical diagnostic (see Fig.
-7), but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [54].
+Individual silhouette widths provide a valuable graphical diagnostic, but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [medoid].
 A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$.
-When even that maximum is low, the limitation often stems from the intrinsic structure of the dissimilarity data rather than a suboptimal choice of algorithm.
+When even that maximum is low, it is possible that the limitation stems from the intrinsic structure of the dissimilarity data.
 This raises the natural question: given a dissimilarity matrix, what is the highest possible $\mathrm{ASW}$?
@@
 is called a silhouette-optimal clustering.
-Once again, recall the toy data (Tab.
-3) and its dissimilarity matrix $\Delta$.
-We evaluate $\operatorname{ASW}(\mathcal{C}_{K},\Delta)$ for every admissible clustering with $K\in\{2,3,4,5\}$ (all set partitions except the trivial K=1K{=}1 case).
-Among the 51 candidates  [^3] The fifth Bell number $B_{5}=52$ and excluding K=1K{=}1 gives $51$., the maximizer is unique:
-$\mathcal{C}^{\ast}\;=\;\bigl{\{}\{x_{1},x_{2},x_{3}\},\;\{x_{4},x_{5}\}\bigr{\}}\quad\text{with}\quad\operatorname{ASW}(\mathcal{C}^{\ast},\Delta)\;=\;0.7512,$
-which matches the per-point silhouettes reported in Tab. 3 and Fig. 7. Finally, the best achievable $\operatorname{ASW}$ within each fixed $K$ is depicted in Tab. 4.
-Table 4: Optimal $\operatorname{ASW}$ within each $K$ on the toy dataset.
-$K$ Best partition Best ASW
-2 $\{x_{1},x_{2},x_{3}\}\mid\{x_{4},x_{5}\}$ 0.7512
-3 $\{x_{1},x_{3}\}\mid\{x_{2}\}\mid\{x_{4},x_{5}\}$ 0.5173
-4 $\{x_{1}\}\mid\{x_{2}\}\mid\{x_{3}\}\mid\{x_{4},x_{5}\}$ 0.3068
-5 $\{x_{1}\}\mid\{x_{2}\}\mid\{x_{3}\}\mid\{x_{4}\}\mid\{x_{5}\}$ 0.0000
-4 The proposed upper bound of the silhouette
-4.1 Overview
+4 A data-dependent upper bound of the silhouette
 We propose a data-dependent ceiling on the ASW that can be computed before any clustering is performed.
 For each data point, we derive a sharp upper bound of its silhouette width.
-Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and moreover, there exists a clustering in which the bound is exactly achieved.
+Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and there exists a clustering in which the bound is exactly achieved.
 Averaging these per-point ceilings produces a single, dataset-level upper bound on the ASW.
-The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and supports optional minimum-cluster-size constraints to make it even sharper.
-4.2 Derivation
+The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and the proposed algorithm supports optional minimum cluster size constraints to further tighten the bound.
 Consider an arbitrary clustering $\mathcal{C}_{K}$ of the dataset $\{1,\ldots,n\}$ with dissimilarites $\Delta$.
@@
 For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$.
-The average distance between $i$ and the other data points in the same cluster is not smaller than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$.
+The average distance between $i$ and the remaining data points in the same cluster is not less than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$.
 Similarly, the average distance between $i$ and the $n-|C_{I}|$ points farthest from $i$ is not smaller than the average distance between $i$ and the data points in the neighboring cluster closest to $i$.
-These two observations are combined to construct a value guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$.
-We formalize this reasoning via the $\Lambda$-quotient of point $i$.
+These two observations are combined to construct a value that is guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$.
+We formalize this reasoning through the $\Lambda$-quotient of point $i$.
 Definition 5.
@@
 $f_{\kappa}(i,\Delta)\coloneqq\min_{\kappa\leq\Lambda\leq n-\kappa}\{q(i,\Delta,\Lambda)\}$
-for some $\kappa\in\{2,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$.
+for some $\kappa\in\{1,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$.
 Often, as we shall see, this leads to strictly tighter limits.
@@
 Algorithm 1 Silhouette upper bound (pointwise)
-The workflow of the proposed upper bound on the ASW is shown in Fig.8.
+The workflow for computing the proposed upper bound is illustrated in Fig.4.
 Following this workflow, Algo.
-1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the (restricted) solution space $\min_{I}|C_{I}|\geq\kappa$.
-The steps are as follows: (1) For each row $i$ of $\Delta$, we sort the off-diagonal entries increasingly to obtain $\hat{\Delta}_{i,1},\hat{\Delta}_{i,2},\ldots$, the nearest-to-farthest distances from $i$.
-(2) For $i$, we initialize the running sums $x$ and $y$ so that $x/(\kappa-1)$ is its mean distance to its $\kappa-1$ nearest points, and $y/(n-\kappa)$ is its mean distance to its $n-\kappa$ farthest points.
-(3) Starting from $\Lambda=\kappa+1$, we update these running sums as $\Lambda$ increases, thereby scanning all admissible values $\Lambda\in\{\kappa,\ldots,n-\kappa\}$.
-For each $\Lambda$, we form the $\Lambda$-quotient $q(i,\Delta,\Lambda)$ and keep the minimum.
-(4) Finally, we record 1q1-q as the upper bound for $s(i\,|\,\mathcal{C}_{K},\Delta)$.
-The array $A$ collects these values across all points, and its average equals the dataset-level upper bound in Eq.
-(1) when $\kappa=1$.
-Complexity
-Forming $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$.
+1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the solution space $\min_{I}|C_{I}|\geq\kappa$.
+The array $A$ collects these values across all points, and its average equals the dataset-level upper bound.
+Time complexity
+The formation of $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$.
 The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall run time is $\mathcal{O}(n^{2}\log n)$.
-This is slightly slower than computing the $\mathrm{ASW}$; $\mathcal{O}(n^{2})$ [54].
+This is slightly slower than computing the $\mathrm{ASW}$, which takes $\mathcal{O}(n^{2})$ time [medoid].
 Output summaries
@@
 (1).
-To reflect the spread of the individual bounds, we denote by $\texttt{minUB}(\Delta)$ and $\texttt{maxUB}(\Delta)$ the minimum and maximum values of $A$, respectively.
+We refer to this as the global upper bound.
 Choosing $\kappa>1$ imposes the restriction that no cluster is allowed to contain fewer than $\kappa$ objects.
 Let $\texttt{UB}_{\kappa}(\Delta)$ denote the $\mathrm{ASW}$ upper bound for this constrained solution space.
+We will refer to $\texttt{UB}_{\kappa}(\Delta)$ as an adjusted upper bound.
 Remark
-The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$; its tightness is data-dependent.
-Nonetheless, it provides an enlightening piece of information that merits inclusion in analyses.
-4.4 Example
-Recall the toy data from Tab.
-3.
-We have $n=5$, so $\Lambda\in\{1,2,3,4\}$.
-The values $q(i,\Delta,\Lambda)$ for all instances are shown in Tab.
-5.
-Consider the K=2K{=}2 clustering from Fig.
-5: $C_{1}=\{x_{1},x_{2},x_{3}\}$ and $C_{2}=\{x_{4},x_{5}\}$.
-Tab.
-6 compares each silhouette width with the corresponding upper bound, illustrating that $s(i)\leq 1-f(i,\Delta)$ for all $i$, with equality for $x_{2},x_{4},x_{5}$ and a small strict headroom for $x_{1}$ ($0.026$) and $x_{3}$ ($0.054$).
-Moreover, we have
-$\operatorname{ASW}(\mathcal{C}^{\ast},\Delta)=0.7512\;\leq\;0.7672=1-\frac{1}{n}\sum_{i=1}^{n}f(i,\Delta).$
-Table 5: All possible values of $q(i,\Delta,\Lambda)$ for each point $i$ in the toy dataset.
-Point $q(\cdot,1)$ $q(\cdot,2)$ $q(\cdot,3)$ $q(\cdot,4)$
-$x_{1}$ 1.000 0.184 0.210 0.466
-$x_{2}$ 1.000 0.417 0.348 0.531
-$x_{3}$ 1.000 0.199 0.253 0.502
-$x_{4}$ 1.000 0.220 0.538 0.643
-$x_{5}$ 1.000 0.213 0.568 0.654
-Table 6: Showing $s(i)\leq 1-f(i,\Delta)$ for each point in the toy dataset.
-$i$ $|C_{I}|$ $\Lambda^{\star}=\arg\min_{\Lambda}q(i,\Delta,\Lambda)$ $f(i,\Delta)$ $1-f(i,\Delta)$ $s(i)$
-$x_{1}$ 3 2 0.184 0.816 0.790
-$x_{2}$ 3 3 0.348 0.652 0.652
-$x_{3}$ 3 2 0.199 0.801 0.747
-$x_{4}$ 2 2 0.220 0.780 0.780
-$x_{5}$ 2 2 0.213 0.787 0.787
-Figure 8: Workflow of the proposed upper bound on the $\mathrm{ASW}$.
-4.5 Enabling early stopping of optimization loops through the upper bound
-We now demonstrate how the data–dependent upper bound introduced in Eq.
-(1) can be used to accelerate standard ASW based optimization loops.
-In many practical applications, selecting the number of clusters involves iterating over $K$ and evaluating the output of the clustering algorithm $\mathcal{A}(K)$ in order to maximize the ASW.
-When the dataset is large, each iteration can be computationally demanding, and exhaustive searches over all candidate $K$ quickly become demanding.
-By incorporating the upper bound into this procedure, we can enable early stopping: if the current ASW is provably within a user–specified relative tolerance of the global optimum, the loop halts, saving potentially substantial computation while still guaranteeing near–optimal clustering performance within the search space.
-This type of procedure is described in more detail in Algo.
-2.
-Input: $\Delta\in\mathbb{R}^{n\times n}$;
-  // Dissimilarity matrix
-1
-Input: $\epsilon$;
-  // Relative error tolerance, $0\leq\epsilon\leq 1$
-2
-Input: $\tau$;
-  // Clusterability threshold, $0\leq\tau\leq 1$
-3
-Input: $\mathcal{A}(K),K_{\max}$;
-  // Clustering algorithm, max. no. clusters
-4
-Output: ASW-optimal clustering of $\Delta$ (within search space)
-Compute $\mathrm{UB}(\Delta)$;
-  // Apply Algo. 1
-5 if $\mathrm{UB}(\Delta)\leq\tau$ then // Abort if dataset is insufficiently clusterable
-6 return NULL
-7optimal clustering $\leftarrow$ NULL ;
-8 optimal ASW $\leftarrow$ -1 ;
-9 for $K=2,\ldots,K_{\max}$ do
-[…]
-11      optimal clustering $\leftarrow\mathcal{A}(K)$ ;
-12      optimal ASW $\leftarrow\mathrm{ASW}(\mathcal{A}(K),\Delta)$ ;
-13    if $\frac{\mathrm{UB}(\Delta)-\text{optimal ASW}}{\mathrm{UB}(\Delta)}<\epsilon$ then
-14       return optimal clustering
-15
-16
-return optimal clustering
-Algorithm 2 Early stopping in ASW-based optimization
-Workflow
-The procedure integrates the proposed upper bound into a standard ASW optimization loop.
-At each step, the current best clustering is compared to the global upper bound.
-If the relative gap falls below the user-specified tolerance $\epsilon$, the algorithm terminates early, certifying that no further search can improve ASW by more than $\epsilon\%$ of the true global optimum.
-In the worst case, the condition is never met, and the loop reverts to the standard exhaustive search over $K=2,\ldots,K_{\max}$, ensuring no loss in clustering quality.
-Remark
-If minimal cluster size constraints are of interest, the procedure can be modified by replacing the initial call to $\mathrm{UB}(\Delta)$ with the restricted version $\mathrm{UB}_{\kappa}(\Delta)$.
-This enables early stopping under user-specified requirements.
-We also emphasise that clusterability in this context is interpreted solely through the lens of the silhouette coefficient.
-While this provides a useful and practical criterion for our purposes, it is by no means a universal or definitive notion of clusterability—many alternative perspectives exist in the literature.
-Recall the toy dataset summarized in Tab.
-4.
-In Sec.
-4.4, we established that the corresponding upper bound is $\texttt{UB}(\Delta)=0.7672$.
-Let $\widehat{S}_{K}$ denote the empirically observed $\mathrm{ASW}$ value for $K$ clusters.
-For the toy dataset, these values are:
-$\widehat{S}_{2}=0.7512,\quad\widehat{S}_{3}=0.5173,\quad\widehat{S}_{4}=0.3068,\quad\widehat{S}_{5}=0.0000.$
-The performance metric of interest is the relative error, defined as
-$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}_{K}}{\mathrm{ASW}(C^{*},\Delta)},$
-where $C^{*}$ denotes the clustering achieving the global maximum ASW.
-Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
-$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}_{K}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}_{K}}{\texttt{UB}(\Delta)}.$
+The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$. Its tightness is data-dependent.
+Use cases
+In theory, when sharp, the upper bound $\texttt{UB}(\Delta)$ can be used to confirm that an empirical ASW has reached the global optimum.
+Even when loose, it increases the interpretability of empirical measurements.
+In real-world scenarios where the global upper bound appears loose, we speculate that adjusted versions can provide more practical use.
+The individual silhouette width upper bounds generated by Algo.
+1 provide a neat complement to standard silhouette visualizations, where the point-wise silhouette widths are typically displayed next to each other in a bar graph.
+In the Experiments section, we examine these use cases.
+Figure 4: Workflow for computing the proposed $\mathrm{ASW}$ upper bound.
+4.2 Extension to the macro-averaged silhouette
+The standard $\mathrm{ASW}$ can produce misleading results when the clusters vary greatly in size.
+In [john], John Pavlopoulos, Georgios Vardakas, and Aristidis Likas show that the so-called macro-averaged silhouette is more robust to cluster-size imbalance.
+For clustering $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$, the macro-averaged silhouette is defined as
+$S_{macro}(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{K}\sum_{I=1}^{K}\frac{1}{|C_{I}|}\sum_{i\in C_{I}}s(i).$
+In practical scenarios where clusters are imbalanced and small groups also matter, this aggregation strategy is more reasonable.
+In this subsection, we use Proposition 1 to construct an upper bound for this silhouette variant in a constrained solution space.
+Consider an empirical clustering result $\widehat{\mathcal{C}}_{K}=\{\widehat{C}_{I}\}_{I=1}^{K}$.
+We narrow our scope in this context to include all clusterings whose cluster sizes match those of $\widehat{\mathcal{C}}_{K}$.
+This narrower solution space is still generally astronomically large.
+Thus, an efficiently computable data-dependent upper bound is of interest as a relevant benchmark.
+Let $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ be any clustering from our constrained solution space, and let $\texttt{ub}_{i}$ denote the silhouette width upper bound of data point $i$, i.e.
+$s(i|\mathcal{C}_{K},\Delta)\leq\texttt{ub}_{i}$.
+Moreover, let $\{\texttt{ub}_{\sigma(i)}\}_{i=1}^{n}$ denote the upper bounds sorted in ascending order, and let $\{N_{I}\}_{I=1}^{K}$ denote the assumed cluster sizes in descending order.
+With $N_{0}\coloneqq 1$, it follows from the rearrangement inequality that
+$S_{macro}(\mathcal{C}_{K},\Delta)$ $\leq\frac{1}{K}\sum_{I=1}^{K}\sum_{i\in C_{I}}\frac{1}{|C_{I}|}\texttt{ub}_{i}$
+$\leq\frac{1}{K}\sum_{I=1}^{K}\frac{1}{N_{I}}\sum_{i=\sum_{J=0}^{I-1}N_{J}}^{\sum_{J=1}^{I}N_{J}}\texttt{ub}_{\sigma(i)}.$
+The last expression is independent of any particular cluster assignment and therefore provides a valid upper bound on the macro-averaged silhouette across all clusterings within the constrained solution space.
+This formulation is particularly useful in applications where cluster sizes are known or can be reasonably estimated.
+Even in the absence of such assumptions, the resulting bound serves as a sharper and more informative benchmark for evaluating empirical clustering results.
+5 Experiments, results & discussion
+Table 1: Real-world datasets. Datasets used in the experiments and their characteristics. The Metric column indicates the distance metrics used.
+Dataset Title Instances Features Classes Field Metric
+Ceramic Chemical Composition of Ceramic Samples 88 17 2 Physics Manhattan
+Customers Wholesale customers 378 7 2 Business Cosine
+Dermatology Dermatology 366 34 6 Health and Medicine Cosine
+Heart Statlog Statlog (Heart) 231 13 2 Health and Medicine Euclidean
+Optdigits Optical Recognition of Handwritten Digits 2296 64 10 Computer Science Euclidean
+Rna Gene expression cancer RNA-Seq 808 20531 5 Biology Euclidean
+Wdbc Wisconsin Diagnostic Breast Cancer 382 30 2 Health and Medicine Euclidean
+Wine Wine 178 13 3 Chemistry Manhattan
+Table 2: Experimental results on synthetic data. Comparison of clustering performance and the proposed upper bound.
+Dataset $K$ ASW $\texttt{UB}(\Delta)$ Worst-case rel. error
+400-64-5-6 5 $0.249$ $0.376$ $0.339$
+400-64-2-2 2 $0.673$ $0.673$ $0.000$
+400-128-7-3 7 $0.522$ $0.566$ $0.078$
+1000-300-5-2 5 $0.668$ $0.678$ $0.015$
+10000-32-20-2 20 $0.626$ $0.774$ $0.192$
+10000-1024-20-4 20 $0.417$ $0.454$ $0.081$
+We evaluate the ASW upper bound by applying it to a variety of datasets, both synthetic and real-world.
+Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity.
+The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
+An essential part of the evaluation is to establish a nearly tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$.
+To this end, we use PAMSIL clustering to obtain an empirical $\mathrm{ASW}$.
+The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies.
+PAMSIL is a suitable algorithm since its objective is to maximize ASW.
+Given an empirical ASW value, say $\widehat{S}$, the performance metric of interest is the relative error, defined as
+$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)},$
+where $C^{*}$ denotes a clustering that achieves the global maximum.
+Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the global upper bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
+$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}}{\texttt{UB}(\Delta)}.$
 We refer to this quantity as the worst-case relative error.
-Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical $\widehat{S}_{K}$ is guaranteed to be within $\epsilon$ of the true maximum ASW.
-Table 7: Early stopping on the toy dataset.
-$K$ $\widehat{S}_{K}$ $\texttt{UB}(\Delta)$ worst-case relative error
-[…]
-[…]
-[…]
-[…]
-As illustrated by Tab.
-7, applying Algo.
-2 with relative error tolerance $\epsilon=0.05$ for the toy dataset would halt the process already at $K=2$, thereby avoiding unnecessary evaluations for $K=3,4,5$.
-In this way, the upper bound allows substantial savings while still certifying near-optimal performance.
-5 Experiments, results & discussion
-5.1 Experimental datasets
-Table 8: Datasets used in our experiments and their characteristics. The Metric(s) column indicates the distance metrics used in the experiments.
-ID Dataset Name Instances Features Classes Field Metric(s)
-1 gene Gene expressions 801 20531 Unknown Genomics Euclidean, correlation
-2 ceramic Ceramic samples 88 17 Unknown Archaeology Euclidean
-3 relig. texts Religious texts 589 8266 Unknown Humanities Cosine, Euclidean, Jaccard
-4 conf. papers Conference papers 5804 11463 Unknown Computer science Cosine, Euclidean, Jaccard
-5 wdbc Breast Cancer Wisconsin Diagnostic 569 30 2 Healthcare Euclidean
-6 wisc Breast Cancer Wisconsin Original 699 9 2 Healthcare Euclidean
-7 ecoli Ecoli 336 7 8 Biology Euclidean
-8 wine Wine 178 13 3 Chemistry Euclidean
-9 iono Ionosphere 351 34 2 Physics Euclidean
-10 zoo Zoo 101 16 7 Biology Euclidean
-11 iris Iris 150 4 3 Botany Euclidean
-12 segment Image Segmentation 2310 19 7 Computer science Euclidean
-13 dermatology Dermatology 366 34 6 Healthcare Euclidean
-14 balance-scale Balance Scale 625 4 3 Psychology Euclidean
-15 heart-statlog Statlog (Heart) 270 13 2 Healthcare Euclidean
-16 thy New Thyroid 215 5 3 Healthcare Euclidean
-We evaluate the upper bound by applying it to a variety of datasets, both synthetic and real-life.
-Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity.
-(For some parts of the experiments we use the Jaccard distance, which means the data is treated in binary form and similarity is based on the amount of shared attributes.)
-An essential part of the evaluation is to establish a tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$.
-To this end, we adopt an empirical approach: we apply several clustering algorithms to each dataset and take the largest observed $\mathrm{ASW}$ as the lower bound.
-The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies.
-The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
-Figure 9: Average silhouette width $\mathrm{ASW}$ vs.
-number of clusters ($K$).
-Results on the synthetic dataset 400-64-5-6 for $K$-means, PAM (PAMSIL), and Hierarchical clustering with weighted and single linkage.
-For each method and each chosen $K$, we compute the resulting $\mathrm{ASW}$.
-The black dashed line marks the theoretical upper bound $\mathrm{UB}(\Delta)=0.376$; the red dashed line marks the upper bound when a minimum cluster size of $80$ is enforced ($0.282$).
-The maximum observed $\mathrm{ASW}$ occurs at K=5K=5—a partition into five equal-sized clusters of 80 elements—and is achieved by $K$-means, Hierarchical (weighted linkage), and PAMSIL.
+Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical ASW is guaranteed to be within $\epsilon$ of the true maximum.
 Synthetic datasets
-We follow [3] in notation and approach.
-Synthetic datasets are generated using the make_blobs function from scikit-learn [68].
+We follow [gaussian] in notation and approach.
+Synthetic datasets are generated using the make_blobs function from scikit-learn [scikit].
 Each dataset is labeled as n_samplesn_featurescenterscluster_stdn\_samples-n\_features-centers-cluster\_std, indicating the parameter values used to generate the points.
 For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion.
-Real-life datasets
-We use sixteen datasets spanning unlabeled (IDs 1–4) and labeled (IDs 5–16) settings (Tab.
-8).
-The unlabeled datasets are named gene expressions, ceramic samples, religious texts, and conference papers.
-For these, we search over $K$ for different algorithms to maximize $\operatorname{ASW}$.
-The labeled datasets consist of twelve standard benchmarks across multiple domains-wdbc, wisc, ecoli, wine, iono, zoo, iris, segment, dermatology, balance-scale, heart-statlog, thy-ranging from 150 to 2,310 instances, 4 to 34 features, and 2 to 8 classes.
-For these datasets, we use the ground-truth class count to set $K$ (e.g., $K=8$ for ecoli), run $K$-medoids to obtain a $K$-clustering, and compare its $\operatorname{ASW}$ with the bound; we also repeat a search over varying $K$ to maximize $\operatorname{ASW}$, mirroring the unlabeled protocol.
-Table 9: Experimental results on synthetic data.
-Each dataset is denoted as n_samples-n_features-centers-cluster_std.
-For Hierarchical clustering (weighted and single linkage), $K$-means, and $K$-medoids (Hybrid), we report the average silhouette width (ASW); the selected number of clusters $K$ is shown in parentheses.
-For $K$-medoids, $\dagger$ indicates that PAMSIL was used; datasets with $\geq\!1000$ samples used PAMMEDSIL to reduce runtime.
-The last three columns summarize the output of Algo.
-1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ upper bound), $\min\mathrm{UB}(\Delta)$, and $\max\mathrm{UB}(\Delta)$ (smallest and largest individual silhouette width upper bound respectively).
-All results use Euclidean distance.
-Dataset Metric
-Weighted
-$K$-means
-$\texttt{UB}(\Delta)$ $\texttt{minUB}(\Delta)$ $\texttt{maxUB}(\Delta)$
-400-64-5-6 Euclidean $0.249$ (5) $0.209$ (4) $0.249$ (5) $0.249^{\dagger}$ (5) $0.376$ $0.276$ $0.467$
-400-64-2-2 Euclidean $0.673$ (2) $0.673$ (2) $0.673$ (2) $0.673^{\dagger}$ (2) $0.673$ $0.619$ $0.718$
-400-128-7-3 Euclidean $0.522$ (7) $0.522$ (7) $0.522$ (7) $0.522^{\dagger}$ (7) $0.566$ $0.514$ $0.619$
-1000-161-2-13 Euclidean $0.084$ (2) $0.068$ (2) $0.084$ (2) $0.046$ (3) $0.182$ $0.137$ $0.256$
-1000-300-5-2 Euclidean $0.668$ (5) $0.668$ (5) $0.668$ (5) $0.668$ (5) $0.678$ $0.651$ $0.706$
-10000-32-20-2 Euclidean $0.626$ (20) $0.626$ (20) $0.626$ (20) $0.626$ (20) $0.774$ $0.681$ $0.847$
-10000-1024-20-4 Euclidean $0.417$ (20) $0.417$ (20) $0.417$ (20) $0.417$ (20) $0.454$ $0.425$ $0.488$
-Table 10: Experimental results on real unlabeled data.
-For each dataset (the distance/dissimilarity used is shown in the “Metric” column), we report the average silhouette width (ASW) for Hierarchical clustering (weighted and single linkage), $K$-means, and $K$-medoids (Hybrid); the selected number of clusters $K$ is given in parentheses.
-For $K$-medoids, $\dagger$ indicates PAMSIL; datasets with $\geq\!1000$ samples used PAMMEDSIL to reduce runtime.
-When the metric is incompatible with $K$-means (e.g., correlation/cosine or Jaccard), the entry is marked N/A.
-The last three columns summarize the output of Algo.
-1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ upper bound), $\min\mathrm{UB}(\Delta)$, and $\max\mathrm{UB}(\Delta)$ (smallest and largest individual silhouette width upper bound respectively).
-Dataset Metric
-Weighted
-$K$-means
-$\texttt{UB}(\Delta)$ $\texttt{minUB}(\Delta)$ $\texttt{maxUB}(\Delta)$
-Gene expressions Correlation $0.345$ (10) $0.332$ (2) N/A $0.395^{\dagger}$ (7) $0.657$ $0.325$ $0.865$
-Religious texts Cosine $0.082$ (234) $0.049$ (2) N/A $0.087^{\dagger}$ (23) $0.349$ $0.079$ $0.810$
-Conference papers Cosine $0.188$ (2) $0.162$ (2) N/A $0.126$ (2) $0.473$ $0.118$ $1.0$
-Religious texts Euclidean $0.846$ (2) $0.846$ (2) $0.428$ (3) $0.846^{\dagger}$ (2) $0.846$ $0.106$ $0.895$
-Ceramic samples Euclidean $0.547$ (2) $0.591$ (2) $0.584$ (2) $0.584^{\dagger}$ (2) $0.855$ $0.561$ $0.950$
-Conference papers Euclidean $0.418$ (2) $0.418$ (2) $0.067$ (2) $0.384$ (2) $0.439$ $0.227$ $1.0$
-Gene expressions Euclidean $0.207$ (2) $0.188$ (2) $0.228$ (6) $0.230^{\dagger}$ (8) $0.419$ $0.174$ $0.624$
-Religious texts Jaccard $0.031$ (106) $0.017$ (569) N/A $0.024^{\dagger}$ (24) $0.119$ $0.040$ $0.719$
-Conference papers Jaccard $0.160$ (2) $0.160$ (2) N/A $0.160$ (2) $0.170$ $0.038$ $1.0$
-Table 11: Early stopping on real data. Early stopping results (Algo. 2) with Euclidean dissimilarity. Best ASW per row is bold; corresponding $K$ is bold.
-Dataset
-$K$
-cand.
-$\epsilon$
-Best ASW
-($K$-means)
-Best ASW
-($K$-medoids)
-UB
-Best $K$
-($K$-means)
-Best $K$
-($K$-medoids)
-Worst-case
-relative error
-($K$-means)
-Worst-case
-relative error
-($K$-medoids)
-Early stop
-($K$-means)
-Early stop
-($K$-medoids)
-gene $[2,15]$ 0.15 0.228 0.230 0.419 6 8 0.46 0.45 false false
-ceramic $[2,30]$ 0.35 0.584 0.584 0.855 2 2 0.32 0.32 true true
-relig. texts $[2,15]$ 0.15 0.428 0.846 0.846 3 2 0.49 0.0 false true
-conf. papers $[2,15]$ 0.15 0.067 0.384 0.439 2 2 0.85 0.13 false true
-wdbc $[2,15]$ 0.15 0.345 0.661 0.708 2 2 0.51 0.07 false true
-wisc $[2,15]$ 0.15 0.574 0.574 0.844 2 2 0.32 0.32 false false
-ecoli $[2,15]$ 0.15 0.398 0.836 0.854 4 2 0.53 0.02 false true
-wine $[2,30]$ 0.35 0.285 0.295 0.626 3 5 0.55 0.53 false false
-iono $[2,15]$ 0.15 0.293 0.413 0.691 6 2 0.58 0.40 false false
-zoo $[2,30]$ 0.35 0.547 0.547 0.839 30 22 0.35 0.35 true true
-iris $[2,30]$ 0.35 0.580 0.580 0.878 2 2 0.34 0.34 true true
-segment $[2,15]$ 0.15 0.357 0.778 0.929 7 2 0.62 0.16 false false
-dermatology $[2,15]$ 0.15 0.277 0.357 0.609 3 2 0.55 0.42 false false
-balance-scale $[2,15]$ 0.15 0.219 0.214 0.737 8 8 0.70 0.71 false false
-heart-statlog $[2,30]$ 0.35 0.169 0.322 0.593 3 2 0.72 0.46 false false
-thy $[2,30]$ 0.35 0.577 0.734 0.835 2 2 0.31 0.12 true true
-5.2 Benchmarking algorithms
-To cluster the datasets, we employ three popular methods: $K$-means, $K$-medoids, and Hierarchical agglomerative clustering.
-These methods serve as baselines to contextualize the performance of the proposed upper bound.
-Because our goal is to obtain the highest possible $\mathrm{ASW}$, we evaluate multiple values of $K$ and report the one that yields the largest $\mathrm{ASW}$.
-$K$-means clustering
-We use the implementation of $K$-means from scikit-learn. For each dataset, the algorithm requires the number of clusters $K$.
-$K$-medoids clustering
-We use the Python package $K$-medoids [77] and, in particular, two medoid-based algorithms: PAMSIL and PAMMEDSIL, both introduced in [52].
-PAMSIL aims to directly optimize $\mathrm{ASW}$ but is computationally expensive.
-PAMMEDSIL is a faster proxy that optimizes a medoid-based version of the silhouette; while expedient, it generally results in a lower $\mathrm{ASW}$ than PAMSIL.
-To balance accuracy and run time, we use PAMSIL for datasets with $n<1000$ and switch to PAMMEDSIL for $n\geq 1000$.
-This threshold was chosen empirically to keep computation tractable while preserving $\mathrm{ASW}$ quality.
-Hierarchical agglomerative clustering (HAC)
-For hierarchical clustering, we use the SciPy implementation [84]. To measure inter-cluster distances when forming the hierarchy, we consider both single and weighted linkage.
-Figure 10: Silhouette plots for the 4 real unlabeled datasets.
-$K$-means clusters; ASW compared against $\texttt{UB}(\Delta)$ and the size-constrained bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size.
-Figure 11: Silhouette plots for the 12 real labeled datasets. $K$-medoids with $K$ fixed to the ground truth; ASW compared against $\texttt{UB}(\Delta)$.
-Figure 12: Silhouette plots for the 12 real labeled datasets.
-$K$-medoids with $K$ chosen by unrestricted optimization; ASW compared with $\texttt{UB}(\Delta)$, illustrating the best empirical quality without fixing $K$ to ground truth.
-Figure 13: Unlabeled datasets: violin plots of pointwise silhouette upper bounds versus minimum cluster size $\kappa$ (up to $\sim 10\%$ of $n$).
-The violins summarize distributions across data points (means, i.e.
-$\mathrm{ASW}$ upper bounds, in red).
-The upper bounds decline as $\kappa$ increases, showing that size constraints lead to sharper limits.
-5.3 Results and discussion
+Finally, for these datasets, we use Euclidean distance.
+Real-world datasets
+We evaluate our method on eight real-world datasets from the UCI Machine Learning Repository [ucidata], commonly used in clustering and classification research.
+Table 1 summarizes their key characteristics along with the distance metrics employed.
+Standard preprocessing steps, including feature scaling and outlier removal, are applied where appropriate.
+Each dataset provides a known number of classes, which we use to determine the number of clusters $K$.
+When all class labels are available, we report the ARI and AMI to verify that these datasets represent meaningful clustering structures.
+In that way, we ensure that our experiments are conducted in natural clustering contexts.
+Table 3: Experimental results on real data. Comparison of clustering performance and the proposed upper bound.
+Performance of PAMSIL clustering
+Comparison with
+global upper bound
+Comparison with
+adjusted upper bound
+Dataset $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI ASW $\texttt{UB}(\Delta)$ Worst-case rel. error $\texttt{UB}_{\kappa}(\Delta)$ Worst-case rel. error
+Ceramic 2 44 $1.000$ $1.000$ $0.358$ $0.621$ $0.423$ $0.403$ $0.112$
+Customers 2 170 $0.394$ $0.366$ $0.400$ $0.951$ $0.579$ $0.574$ $0.303$
+Dermatology 6 21 $0.850$ $0.875$ $0.455$ $0.836$ $0.456$ $0.704$ $0.354$
+Heart Statlog 2 83 $0.447$ $0.349$ $0.161$ $0.600$ $0.731$ $0.279$ $0.422$
+Optdigits 10 50 $0.803$ $0.819$ $0.180$ $0.620$ $0.709$ $0.463$ $0.611$
+Rna 5 77 $0.225$ $0.419$ $0.464$ $0.323$ $0.303$
+Wdbc 2 46 $0.427$ $0.386$ $0.348$ $0.621$ $0.440$ $0.455$ $0.236$
+Wine 3 53 $0.884$ $0.864$ $0.315$ $0.649$ $0.514$ $0.450$ $0.298$
+Figure 5: Silhouette plots for the eight real-world datasets.
+PAMSIL clustering with $K$ equal to the ground truth number of classes; ASW compared against the global upper $\texttt{UB}(\Delta)$ and the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size.
+The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds.
+Figure 6: PAMSIL performance for varying $K$ and comparison with upper bounds.
+Results
+In Table 2, we apply the upper bound to six synthetic datasets and compare the results with PAMSIL clustering.
+For dataset 400-64-2-2, the upper bound at $0.673$ confirms that PAMSIL reaches a globally optimal silhouette score.
+For three of the remaining datasets, $\texttt{UB}(\Delta)$ proves that the PAMSIL ASW is within 8 $\%$ of the global maximum, that is, essentially optimal.
+For dataset 400-64-5-6 with $\texttt{UB}(\Delta)=0.376$, we record an empirical ASW of $0.249$.
+In this case, it remains unclear whether further substantial improvements of the ASW are possible.
+However, interpreting the attained value in the interval [1,0.376][-1,0.376] enables more well-founded conclusions compared to using the standard range [1,1][-1,1].
+Table 3 examines the performance of PAMSIL on the eight real-world datasets and compares it with the upper bounds.
+In these cases, the global upper bounds provide information that is less actionable.
+It is unclear whether the reason for this is that the attained ASW values are far from $\mathrm{ASW}(C^{*},\Delta)$, or that the upper bounds are somewhat loose, or both.
+However, the adjusted upper bounds $\texttt{UB}_{\kappa}(\Delta)$ are considerably tighter for most datasets.
+Here, $\kappa$ is set to the minimum cluster size in the PAMSIL solution.
+For datasets Ceramic, Customers, Rna, Wdbc and Wine, $\texttt{UB}_{\kappa}(\Delta)$ proves that the ASW-values attained by PAMSIL are within 30 $\%$ of the corresponding optimums.
 Fig.
-9 presents the $\mathrm{ASW}$ obtained by different clustering algorithms on the synthetic dataset 400-64-5-6.
-K-means achieves its best result with five clusters of equal size ($80$ objects each), reaching an $\mathrm{ASW}$ of $0.249$.
-Hierarchical clustering with weighted linkage also favors five clusters, whereas single linkage attains its maximum $\mathrm{ASW}$ ($0.209$) at four clusters.
-K-medoids reaches the same output as K-means and weighted linkage.
-The theoretical upper bound without restrictions on cluster size ($\kappa=1$) is $0.376$ (black dashed line).
-When enforcing a minimum cluster size of $\kappa=80$, Algo.
-1 yields a reduced upper bound of $0.282$ (red dashed line).
-Thus, among solutions with no cluster smaller than $80$ elements, the five-cluster partition found by e.g.
-K-means is guaranteed to be within $12\%$ of the global optimum.
-Tab.
-5.1 reports the remaining results on the synthetic datasets.
-For each algorithm, we show the best observed $\mathrm{ASW}$, with the selected number of clusters in parentheses.
-The last three columns summarize the upper bounds from Algo.
-1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ bound), $\texttt{minUB}(\Delta)$ (smallest individual bound), and $\texttt{maxUB}(\Delta)$ (largest individual bound).
-For dataset 400-64-2-2, the upper bound matches the achieved $\mathrm{ASW}$ for all algorithms, certifying global optimality.
-Dataset 1000-161-2-13 has a significantly low upper bound ($\mathrm{UB}(\Delta)=0.182$), implying that no clustering can exceed an $\mathrm{ASW}$ of 0.182; in such cases, alternative quality measures may be more appropriate, or the data may be effectively non-clusterable.
-Tab.
-5.1 reports results on the real unlabeled datasets and mirrors the layout of Tab.
-5.1.
-By default, we search over $K\in\{2,\ldots,n\}$, where $n$ is the number of samples.
-When necessary for reasonable run time, however, we limit the search at $K\leq 30$.
-Entries marked N/A indicate that $K$-means is incompatible with the metric, e.g., correlation, cosine or Jaccard.
-For Euclidean “Religious texts”, Euclidean “Conference papers”, and Jaccard “Conference papers”, the hierarchical methods are within $0.021$ of the upper bound $\mathrm{UB}(\Delta)$—so any further improvement in ASW over the reported values is at most $0.021$.
-Several datasets have small $\texttt{minUB}(\Delta)$ (e.g., Cosine “Religious texts” with $\texttt{minUB}(\Delta)=0.079$), certifying that at least one observation cannot achieve a silhouette width exceeding that value under any clustering.
-In general, when $\texttt{minUB}(\Delta)$ is near zero, it may be reasonable to treat the corresponding observations as noise or remove them from the analysis.
-We implement the early-stopping algorithm presented in Sec.
-4.5 and examine empirically how it performs on the real datasets.
-In Tab.
-11, we adopt the following convention: For datasets with $n\leq 300$, we explore $K\in\{2,\dots,30\}$ with tolerance $\epsilon=0.35$.
-For datasets with $n>300$, we restrict the search to $K\in\{2,\dots,15\}$ with tolerance $\epsilon=0.15$.
-These parameter choices are not canonical, but rather serve to illustrate different practical regimes.
-The goal is to demonstrate how Algo.
-2 behaves under varying $K$ candidates and tolerance, rather than to prescribe a universally optimal setting.
-The results indicate that the proposed upper bound can indeed shorten silhouette-based optimization loops in practice under reasonable tolerance choices.
-In several cases — notably ceramic, religious-texts, conference-papers, wdbc, ecoli, zoo, iris, and thy — the algorithm halted early while still returning an ASW within the specified tolerance.
-In other cases, the bound was not tight enough compared to the observed output to trigger early stopping, and the procedure effectively reverted to a full search.
-This mixed outcome highlights both the promise and the limitations of the approach: when the bound interacts favorably with the empirical ASW, it can reduce computation substantially, but there is no guarantee of savings in every scenario.
-Fig.
-10 visualizes for the real unlabeled datasets the individual upper bounds in the context of the clusterings obtained by K-means (see Tab.
-5.1).
-It also shows how the upper bound is tightened under size constraints.
-Fig.
-11 and Fig.
-12 are constructed in the same way, but for the labeled datasets, without size constraints and with K-medoids instead of K-means.
-The difference between the two of them is that the latter implements a scan of different $K$ values to maximize the empirical ASW whereas the former uses the ground truth labels to stipulate $K$.
-Fig.
-13 also corresponds to the labeled data and illustrates a more rigorous analysis of the distribution of the individual upper bounds in relation to the $\kappa$ parameter of Algo.
-1.
-We conclude the discussion by highlighting three key roles of the proposed upper bound.
-  1. 1.
-    Global upper bound.
-The unrestricted bound $\mathrm{UB}(\Delta)$ is a ceiling on any achievable ASW for the dataset.
-It can shorten silhouette-based search loops by enabling early stopping; when tight, it certifies that a globally silhouette-optimal clustering has been found.
-Even when not tight, it provides essential context for interpreting reported ASW values across algorithms and initializations.
-Moreover, the magnitude of $\mathrm{UB}(\Delta)$ itself can be seen as a proxy for the clusterability of the dataset: a low upper bound suggests that no clustering can yield well-separated and compact groups, whereas a high upper bound indicates the potential for meaningful structure.
-  2. 2.
-    Restricted upper bound.
-Analysts often care about a constrained solution space.
-By enforcing a minimum cluster size via $\kappa$, the sharper bound $\mathrm{UB}_{\kappa}(\Delta)$ reflects application constraints and yields insights that are more relevant.
-  3. 3.
-    Individual upper bounds.
-Our proposed pointwise upper bounds quantify the maximum silhouette each observation can attain.
-Very low values flag observations that cannot be well integrated under any clustering; such points may merit down-weighting, manual inspection, or exclusion as outliers or noise.
+5 shows classical silhouette visualizations corresponding to the PAMSIL clusters, complemented with upper bounds of the individual silhouette widths.
+The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds.
+For the adjusted upper bounds, $\kappa$ is set to the smallest cluster size in the generated solution.
+This method provides a visual comparison of the empirical silhouette and the highest possible value at the data point level.
+In Fig.
+6 we show, for each real-world dataset and for each $K\in\{2,\ldots,15\}$, the attained ASW value compared to the global ASW upper bound.
+Additionally, at each $K$, we compare the ASW to the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$ with $\kappa$ set to the minimum cluster size in the corresponding solution.
+Lastly, we include, at each $K$, the recorded macro-averaged silhouette and compare this value with the upper bound derived in Section 4.2.
+Discussion
+We summarize with some reflections on the practical relevance of the proposed upper bound.
+Replacing the standard $\mathrm{ASW}$ range [1,1][-1,1] with the data-dependent interval $1,\texttt{UB}(\Delta)]$ enhances the interpretability of observed $\mathrm{ASW}$ values.
+However, when the empirical $\mathrm{ASW}$ lies far below the upper bound, it remains unclear whether further improvement is actually possible.
+We suspect that this gap often persists for the global upper bound, even when the empirical $\mathrm{ASW}$ is close to $\mathrm{ASW}(C^{*},\Delta)$.
+In many real-world applications, however, analysts impose minimum cluster-size constraints to avoid overly small or degenerate clusters.
+By incorporating such constraints through $\kappa$, the refined bound $\texttt{UB}_{\kappa}(\Delta)$ better reflects the relevant solution space and provides more meaningful guidance in practice.
+As demonstrated in Table 3, $\texttt{UB}_{\kappa}(\Delta)$ ensures that the PAMSIL solutions are within 30 $\%$ of the optimum for five of the examined datasets, indicating strong empirical performance under realistic conditions.
+In such cases, the bound can be used to conclude that further exploration would, at best, yield only marginally better results.
 6 Conclusion
-In this paper, we propose a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a new tool for analyzing clustering quality.
-While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a much tighter, data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time.
-This upper bound provides a more meaningful reference value: it often lies well below $1$, reflecting the inherent limitations of the dataset itself.
-Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often close-to-sharp and yields valuable insights into the structure and limitations of the data.
+In this paper, we proposed a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a complementary tool for analyzing clustering quality.
+While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a tighter data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time.
+This upper bound provides a more meaningful reference value.
+It often lies well below $1$, reflecting the inherent limitations of the dataset itself.
+Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often nearly tight and yields valuable insights into the structure and limitations of the data.
 At the same time, we emphasize that the bound is not guaranteed to be sharp in general; the true global ASW-maximum may fall considerably below the bound.
-For this reason, the bound should be viewed not as a precise target but as an informative ceiling that frames observed ASW scores in a broader context.
-Even when loose, the bound can shorten optimization loops, offer a first-order estimate of the datasets clusterability, and highlight points that are inherently difficult to cluster.
+For this reason, the bound should not be viewed as a precise target but as an informative ceiling that frames observed ASW scores in a sharper context.
 From a broader perspective, our main contribution is to establish a proof of concept: that efficiently computable, data-dependent upper bounds on ASW can enrich cluster analysis.
-We believe this idea opens a promising line of work, where future research should aim at deriving tighter bounds, extending the approach to other clustering validity indices, and exploring additional constraints motivated by application domains.
-A particularly important direction is to develop theoretical guarantees, for example bounding how far the proposed upper bound can deviate from the true global maximum under certain assumptions on the dataset.
-This goal could be attained through the derivation of a theoretical and close-to-sharp lower bound on $\mathrm{ASW}(\mathcal{C}^{*},\Delta)$.
+We believe this idea opens a promising direction for future work, where the focus should be on deriving tighter bounds and extending the framework to other internal clustering validity indices.
+An important next step is to better understand under which conditions the proposed bound is most informative in practice — for instance, how its tightness varies with properties of the data, the dissimilarity metric, the dimensionality of the feature space, or the number of underlying clusters.
+Such insights would clarify in which contexts the bound provides the greatest practical value.
 Declaration of competing interest
@@
 Acknowledgments
-The authors thank Hugo Fernström for his careful proofreading and for offering insightful feedback on a previous version.
-His constant encouragement greatly contributed to the development of this work.
+The authors would like to thank Hugo Fernström for his careful proofreading and thoughtful feedback on an earlier version of this manuscript.
+His continuous encouragement has been greatly appreciated throughout the development of this work.
+We are also grateful to several researchers for their valuable comments and suggestions on a previous version of the paper.

### v4

@@
 Abstract
-The silhouette coefficient quantifies, for each observation, the balance between within-cluster cohesion and between-cluster separation, taking values in [1,1][-1,1].
-The average silhouette width (ASW) is a widely used internal measure of clustering quality, with higher values indicating more cohesive and well-separated clusters.
-However, the dataset-specific maximum of ASW is typically unknown, and the standard upper limit of 1 is rarely attainable.
-In this work, we derive for each data point a sharp upper bound on its silhouette width and aggregate these to obtain a canonical upper bound on the ASW.
-This bound—often substantially below 1—enhances the interpretability of empirical ASW values by indicating how close a given clustering result is to the best possible outcome on that dataset.
-It can be used to confirm global optimality, guide the evaluation of clustering solutions, and be refined to incorporate minimum cluster-size constraints for greater practical relevance.
-Finally, we extend the framework to establish a corresponding bound for the macro-averaged silhouette.
+The silhouette coefficient quantifies, for each observation, the balance between within-cluster cohesion and between-cluster separation, taking values in the range [1,1][-1,1].
+The average silhouette width ($\operatorname{ASW}$) is a widely used internal measure of clustering quality, with higher values indicating more cohesive and well-separated clusters.
+However, the dataset-specific maximum of $\operatorname{ASW}$ is typically unknown, and the standard upper limit of 1 is rarely attainable.
+In this work, we derive for each data point a sharp upper bound on its silhouette width and aggregate these to obtain a canonical upper bound of the $\operatorname{ASW}$.
+This bound—often substantially below 1—enhances the interpretability of empirical $\operatorname{ASW}$ values by providing guidance on how close a given clustering result is to the best possible outcome for that dataset.
+We evaluate the usefulness of the upper bound on a variety of datasets and conclude that it can meaningfully enrich cluster quality evaluation; however, its practical relevance depends on the specific dataset.
+Finally, we extend the framework to establish an upper bound of the macro-averaged silhouette.
 keywords:
@@
 1 Introduction
-Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [aggarwal2013introduction].
-Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [dinh2025data].
-Because ground-truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [jain1999data].
-Among the cluster validation indices, the silhouette width (also known as the silhouette coefficient, silhouette score, or silhouette index) is one of the most widely used measures to assess how well each point fits within its assigned cluster relative to its nearest neighboring cluster [ogsilhouette].
-The average of this measure across all data points is often referred to as the average silhouette width $\mathrm{ASW}$.
-$\mathrm{ASW}$ values close to 1 signal well‑separated, compact clusters; values near 0 imply overlapping clusters; and negative values indicate that points are likely misassigned.
-Extensive work has examined and refined $\mathrm{ASW}$.
-[batool2021clustering] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency.
-More recently, [john] examined alternative aggregation strategies that improve $\mathrm{ASW}$’s robustness to cluster-size imbalance.
-Although $\mathrm{ASW}$ is widely used, its raw values are difficult to interpret across heterogeneous datasets.
-Intrinsic data characteristics impose strict ceilings on attainable $\mathrm{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\mathrm{ASW}$.
-Comparing $\mathrm{ASW}$ values across datasets with different properties can therefore lead to misguided conclusions about cluster quality.
-To address this issue, we derive a data-dependent upper bound on the $\mathrm{ASW}$.
-This bound supplies essential context: if analysis shows that no partition can exceed, say, an $\mathrm{ASW}$ of 0.30, an empirical result of 0.29 should be regarded as essentially optimal.
-Such insight prevents wasted effort in pursuit of negligible gains and enables more informed evaluation of clustering quality.
+Cluster analysis is a fundamental tool in data science for discovering structure in unlabeled data [1].
+Its applications span the natural and social sciences, engineering, education, economics, and the health sciences [9].
+Because ground truth labels are rarely available, evaluating clustering quality is challenging, so researchers rely on internal validation indices that summarize within-cluster compactness and between-cluster separation [11].
+Among the internal cluster validation indices, the silhouette width (also known as the silhouette coefficient, silhouette score, or silhouette index) is one of the most widely used measures.
+It quantifies how well each point fits within its assigned cluster relative to its nearest neighboring cluster [18].
+The average of this measure across all data points is often referred to as the average silhouette width ($\operatorname{ASW}$).
+Silhouette scores close to 1 signal well-separated, compact clusters; values near 0 typically occur for points that lie between two clusters; a negative value indicates that a point is on average closer to its nearest neighboring cluster than to its own and is therefore commonly used to flag a potential misplacement.
+These interpretations are most reliable when clusters are roughly isotropic and of comparable scale; when cluster diameters differ substantially, silhouette values can be misleading, a limitation we return to in Section 6.
+Extensive work has examined and refined $\operatorname{ASW}$.
+Batool et al.
+[4] proved that it satisfies key axioms for clustering quality measures such as scale invariance and consistency.
+More recently, Pavlopoulos et al.
+[16] examined alternative aggregation strategies that improve $\operatorname{ASW}$’s robustness to cluster size imbalance.
+The so-called medoid-based silhouette is a simple approximation of the silhouette coefficient that replaces average distances to clusters with distances to cluster medoids, making computations cheaper [14].
+The medoid of a cluster is the cluster member with the smallest total distance to all other members [20].
+Although $\operatorname{ASW}$ is widely used, its raw values are difficult to interpret.
+Intrinsic data characteristics impose strict ceilings on attainable $\operatorname{ASW}$ scores; for example, when clusters overlap or deviate from convex shapes, even the best partition may yield a low $\operatorname{ASW}$.
+As a result, observing a low $\operatorname{ASW}$ value alone does not reveal whether this reflects poor clustering performance or an inherent limitation of the data itself.
+For any dataset with more than a few samples, the true maximum ASW is unknown, which makes it difficult to assess clustering performance in relative terms.
+Relating an empirical $\mathrm{ASW}$ to a data-dependent upper bound could mitigate this problem by indicating how close the given solution is to the best possible outcome.
+Against this background, we pose the following research question.
+Question: Given a dissimilarity matrix, can we efficiently compute an upper bound of the $\operatorname{ASW}$ that provides meaningful guidance on how close an empirical clustering is to the global $\operatorname{ASW}$ maximum?
+In connection with this question, our main contributions are as follows:
+    We introduce a novel data-dependent upper bound on the $\operatorname{ASW}$, computable in $O(n^{2}\!\log n)$ time, to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass.
+    We evaluate the usefulness of the proposed upper bound on synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures.
+    All datasets, preprocessing scripts, bound computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
+In theory, the proposed upper bound can provide important context for interpretation: if analysis shows that no partition can exceed, say, an $\operatorname{ASW}$ of $0.30$, then an empirical result of $0.29$ can be understood as being close to the best achievable value for the given data.
+Such information helps practitioners assess how much room for improvement remains and avoid expending effort in search of gains that are provably limited by the structure of the data.
 The biggest obstacle in silhouette optimization is the combinatorial explosion of the solution space as the number of samples increases.
 Even for moderately sized datasets, the number of possible clusterings is astronomical, rendering optimization through exhaustive search with respect to any internal criterion computationally prohibitive.
-[batool2021clustering] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\mathrm{ASW}$, noting that global optimization is tractable only for the smallest datasets.
+Batool et al.
+[4] addressed this limitation by proposing two heuristic algorithms that search for local maxima of $\operatorname{ASW}$, noting that global optimization is tractable only for the smallest datasets.
 Our contribution complements their approach by providing a data-aware benchmark against which the quality of such locally optimal solutions can be evaluated.
-Figure 1: Illustration of the silhouette coefficient calculation.
-The main contributions are as follows:
-    We introduce a novel data–dependent upper bound on the $\mathrm{ASW}$, computable in $O(n^{2}\!\log n)$ time, to yield a global ceiling that no clustering of the given dissimilarity matrix can surpass.
-    We evaluate the proposed upper bound on large-scale synthetic datasets and a diverse suite of real-world datasets across multiple dissimilarity measures.
-We show that the bound is nearly tight in many of the cases examined.
-    All datasets, preprocessing scripts, bound computation routines, and experiment notebooks are publicly released in a GitHub repository [^1] https://github.com/hugo-strang/silhouette-upper-bound and PyPI [^2] https://pypi.org/project/silhouette-upper-bound/ to ensure full reproducibility and to encourage further research and practical adoption.
+Figure 1 illustrates the intended workflow.
+1. Compute upper bound2. Measure empirical ASW3. Estimate proximity toglobal optimality
+Figure 1: Illustrative workflow for using the upper bound as a diagnostic tool.
+The process involves (1) computing the $\mathrm{ASW}$ upper bound from the given dissimilarity matrix, (2) measuring the $\mathrm{ASW}$ achieved by the chosen clustering algorithm, and (3) assessing the proximity to global optimality by analyzing the gap between these two values.
+To summarize: this paper’s main objective is to propose and assess a method for computing an upper bound of the $\mathrm{ASW}$ for a given dissimilarity matrix.
+We believe this contribution is important because such a data-dependent upper bound could be useful in clustering quality evaluations.
 The rest of this paper is organized as follows.
@@
 Section 4 details the proposed method.
-Section 5 presents the experimental results.
+Section 5 presents and discusses the experimental results.
 Finally, Section 6 summarizes the findings and outlines directions for future research.
+Lastly, Table 1 summarizes the notation used throughout the paper.
+Table 1: Summary of notation used throughout the paper.
+Symbol Description
+$n$ Number of objects in the dataset
+$i,j$ Indices of objects, in $\{1,\dots,n\}$
+$\Delta\in\mathbb{R}^{n\times n}$ Dissimilarity matrix with entries $\Delta_{ij}$
+$\hat{\Delta}\in\mathbb{R}^{n\times(n-1)}$ Row-wise sorted off-diagonal dissimilarities
+$K$ Number of clusters
+$\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ A $K$-clustering of $\{1,\dots,n\}$
+$C_{I}$ Cluster $I$; $|C_{I}|$ denotes its size
+$\mathcal{C}^{*}$ Silhouette-optimal clustering
+$a(i)$ Average dissimilarity from object $i$ to members of its own cluster
+$b(i)$ Minimum average dissimilarity from $i$ to points in any other cluster
+$s(i\mid\mathcal{C}_{K},\Delta)$ Silhouette width of object $i$
+$\operatorname{ASW}(\mathcal{C}_{K},\Delta)$ Average silhouette width
+$S_{\mathrm{macro}}(\mathcal{C}_{K},\Delta)$ Macro-averaged silhouette
+$q(i,\Delta,k)$ $k$-quotient of object $i$ used in the upper bound construction
+$m$ Minimum allowed cluster size in constrained solution spaces
+$f(i,\Delta)$, $f_{m}(i,\Delta)$ Minima of $q(i,\Delta,k)$ over $k$ (unconstrained and with $\min_{I}|C_{I}|\geq m$)
+$\texttt{UB}(\Delta)$ Global upper bound of $\operatorname{ASW}$
+$\texttt{UB}_{m}(\Delta)$ Upper bound of $\operatorname{ASW}$ with minimum cluster size $m$
 2 Related work
-Silhouette coefficientEnergy & Power Systems[ouyang2024k][rodriguez2024novel][zheng2025practical][duarte2021increasing][eskandarnia2022embedded][zhao2023generalized][koutsandreas2024harnessing][zhou2025recognition][hu2023transformer][alrasheed2024building][diaz2022typologies]Transportation & Mobility[mohd2024passengers][wang2021analyzing][bai2023adaptive][sheng2025multi][guo2025adaptive][passarella2024anomaly][ma2025unravelling][li2025clustering][ansari2021spatiotemporal]Biomedical & Health[jamalpour2025high][sharafoddini2021identifying][ghaderi2023self][gill2024symptom][baier2024nocturnal][kim2022pathological][rustam2022data][jha2021apache][xian2025systematic][bhagat2022dpcf]Food & Agriculture[edris2024identifying][singh2025development][gere2023recommendations]Social & Economic[owoo2025hierarchical][quintero2022analysis][gulma2022new][zhang2025two][aonso2024cluster]Materials & Chemistry[pham2025clustering][bi2025comprehensive][sancho2022cluster][bagirov2023novel][consuegra101unsupervised]Environment & Earth Sciences[nourani2022robust][guo2023combining][yang2022zoning][huang2023dominant][balbi2024hierarchical][biabiany2020design][guan2025novel][gupta2025criteria][zhao2023comparison][deng2025on]Sensor Networks & IoT[kotary2020distributed][kotary2021many][panwar2024distributed][panwar2025distributed][waidyanatha2023regularity]Big Data[qaroush2025maximum][thirumoorthy2021hybrid][song2023novel][chander2023dolphin]Cluster Validity & Selection[khan2025standardizing][khan2025standardization][batool2021clustering][lenssen2024medoid][capo2023fast][gagolewski2021are][bagirov2023finding]Algorithm Development[koren2025preprocessing][silva2025automatic][manochandar2020development][qin2025single][zhang2025two][bai2023adaptive][sheng2025multi][kotary2020distributed][kotary2021many][du2024multi][fu2025gsscl][gan2024discrimination][fu2023gc][enriquez2025performance][gere2023recommendations]Feature & Dim‑Reduction[rustam2022data][dwivedi2025novel][zhou2024ocae][chander2023dolphin]
-Figure 2: Taxonomy of how the silhouette coefficient is employed across research fields.
-The silhouette coefficient [ogsilhouette] is a distance-based internal validity index that quantifies how well each object fits into its assigned cluster relative to its nearest alternative.
-Fig.
-1 illustrates the construction.
-For a point $x_{i}$ in cluster $C_{\ell}$, we measure within-cluster cohesion by $a(x_{i})$—its average dissimilarity to the other members of $C_{\ell}$—and nearest-cluster separation by $b(x_{i})$—its average dissimilarity to the closest neighboring cluster.
+Figure 2: Illustration of the silhouette coefficient calculation.
+The silhouette coefficient [18] is a distance-based internal validity index that quantifies how well each object fits into its assigned cluster relative to its nearest alternative.
+Figure 2 illustrates the construction.
+For a point $x_{i}$ in cluster $C_{I}$, we measure within-cluster cohesion by $a(x_{i})$—its average dissimilarity to the other members of $C_{I}$—and nearest-cluster separation by $b(x_{i})$—its average dissimilarity to the closest neighboring cluster.
 These quantities yield an individual silhouette score $s(x_{i})\in[-1,1]$, which is large when separation dominates cohesion, near zero around decision boundaries, and negative when $x_{i}$ is, on average, closer to another cluster than to its own.
-Aggregating over points gives the ASW, a dataset-level summary of partition quality.
-In practice, ASW is often used for model selection by choosing the number of clusters $K$ that maximizes it.
-Despite numerous extensions—e.g., to categorical data [dinh2019estimating], fuzzy clustering [lasek2024employing], hierarchical clustering [owoo2025hierarchical], model-based clustering [guan2025novel], spectral clustering [xian2025systematic], spherical clustering [deng2025on], and optimization-based methods [yang2024clustering]—the original silhouette framework remains a standard baseline for benchmarking clustering algorithms.
-Other commonly used internal validation indices include the Davies–Bouldin index [dbindex] and the Calinski–Harabasz index [chindex], which also evaluate compactness and separation but rely on alternative formulations of intra- and inter-cluster dispersion.
+Aggregating over points gives the $\operatorname{ASW}$, a dataset-level summary of partition quality.
+In practice, $\operatorname{ASW}$ is often used for model selection by choosing the number of clusters $K$ that maximizes it [2, 25, 3].
+Other commonly used internal validation indices include the Davies–Bouldin index [7] and the Calinski–Harabasz index [5], which also evaluate compactness and separation but rely on alternative formulations of intra- and inter-cluster dispersion.
 In this work, we focus on the silhouette due to its interpretability and widespread adoption across both research and applied domains.
-To account for cluster-size imbalance, we additionally employ the macro-averaged silhouette, defined as the unweighted mean of per-cluster silhouette averages.
+To account for cluster size imbalance, we additionally employ the macro-averaged silhouette [16], defined as the unweighted mean of per-cluster silhouette averages.
 This variant ensures that smaller clusters contribute equally to the overall assessment, providing a more balanced view of clustering quality.
 Silhouette-optimization
-PAMSIL (Partitioning Around Medoids based on SILhouette) is a modification of the classical PAM algorithm [kaufman90], designed to directly optimize the ASW [pamsil].
-Unlike $K$-means or standard $K$-medoids, which minimize within-cluster dissimilarities, PAMSIL explicitly seeks cluster configurations that maximize ASW.
-It achieves this by modifying the SWAP phase of PAM: in each iteration, the algorithm performs the swap that yields the largest improvement in ASW and terminates once no further improvement is possible, ensuring convergence to a local optimum.
-An efficient implementation of PAMSIL is available in the open-source K-Medoids Python package [Schubert2022].
-Moreover, Fatima Batool and Christian Hennig have proposed alternative algorithms that find local ASW optima [batool2021clustering].
+PAMSIL (Partitioning Around Medoids based on SILhouette) is a modification of the classical PAM algorithm [12], designed to directly optimize the $\operatorname{ASW}$ [22].
+Unlike $K$-means or standard $K$-medoids, which minimize within-cluster dissimilarities, PAMSIL explicitly seeks cluster configurations that maximize $\operatorname{ASW}$.
+It achieves this by modifying the SWAP phase of PAM: in each iteration, the algorithm performs the swap that yields the largest improvement in $\operatorname{ASW}$ and terminates once no further improvement is possible, ensuring convergence to a local optimum.
+An efficient implementation of PAMSIL is available in the open-source kmedoids Python package [19].
+Moreover, Fatima Batool and Christian Hennig have proposed alternative algorithms that find local $\operatorname{ASW}$ optima [4].
+In practice, PAMSIL can be computationally demanding, particularly for larger datasets.
+For this reason, accelerated variants have been developed.
+The kmedoids package includes FastMSC and DynMSC, which are both methods for optimizing the medoid-based version of the silhouette.
+DynMSC automatically explores different values of $K$ and returns the clustering with the highest average medoid silhouette.
+Despite these developments around the silhouette coefficient, the existing literature largely focuses on 1) comparing the silhouette score with other clustering quality indices [6], 2) proposing closely related alternatives that are preferable in certain contexts [16, 22], or 3) designing algorithms for silhouette-optimization [22, 4, 14, 19].
+To the best of our knowledge, these works do not provide an efficiently computable, data-dependent upper bound on the best achievable $\mathrm{ASW}$ for a given dataset.
+Our contribution complements these lines of work by deriving such a bound directly from the pairwise dissimilarities, enabling empirical silhouette scores to be interpreted relative to a dataset-specific ceiling and providing a principled benchmark for silhouette-optimization heuristics.
 External validation indices
-When ground-truth class labels are available, external validation measures provide a way to quantify how well a clustering agrees with a known reference partition.
-Two of the most commonly used indices are the adjusted Rand index (ARI) [ari] and the adjusted mutual information (AMI) [ami].
+When ground truth class labels are available, external validation measures provide a way to quantify how well a clustering agrees with a known reference partition.
+Two of the most commonly used indices are the adjusted Rand index (ARI) [10] and the adjusted mutual information (AMI) [23].
 Both adjust their underlying similarity measures for chance agreement, ensuring that random labelings yield expected scores near zero.
-The ARI extends the classical Rand index by comparing all pairs of samples and counting how consistently they are assigned to the same or different clusters in both partitions.
-Its normalization accounts for random overlap, producing values in [1,1][-1,1] with $1$ indicating identical partitions.
+The ARI extends the classical Rand index, which compares all pairs of samples and counts how consistently they are assigned to the same or different clusters in both partitions.
+Its normalization accounts for random overlap and produces values in [1,1][-1,1] with $1$ indicating identical partitions.
 The AMI applies a similar correction to the mutual information (MI) between partitions, normalizing it by the expected MI under random assignments and bounding the result between 0 and $1$.
-Both indices are symmetric, easy to interpret, and widely used in benchmarking studies, where they enable consistent comparison of clustering algorithms across datasets with known class structure.
+Both indices are symmetric and widely used in benchmarking studies, where they enable consistent comparison of clustering algorithms across datasets with known class structure.
 3 Preliminaries
@@
 The triangle inequality is not required. We also exclude matrices containing a row of all zeros.
-Figure 3: The first 27 Bell numbers, which enumerate the partitions of an $n$-element set. This rapid growth highlights the combinatorial explosion of clustering possibilities as $n$ increases.
-For our purposes, it will be convenient to work with the off-diagonal entries of $\Delta$, sorted row-wise in ascending order.
-We denote this transformed dissimilarity matrix by $\hat{\Delta}$ ($\hat{\Delta}\in\mathbb{R}^{n\times n-1}$).
-In other words, we have for every $i$ $\{\Delta_{ij}\}_{j=1}^{n}\setminus\{\Delta_{ii}\}=\{\hat{\Delta}_{ij}\}_{j=1}^{n-1}$ and $\hat{\Delta}_{i1}\leq\hat{\Delta}_{i2}\leq\cdots\leq\hat{\Delta}_{i(n-1)}$.
-Constructing $\hat{\Delta}$ requires sorting each of the $n$ rows of $\Delta$, which costs $\mathcal{O}(n^{2}\log n)$ time in total.
 Now, let us formally define the concept of a clustering of the dataset $\{1,\ldots,n\}$.
 Definition 2 (Clustering).
-Let $K\in\{2,\ldots,n\}$. A $\mathbf{K}$-clustering of $\{i\}_{i=1}^{n}$ is a set $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ such that $C_{I}\subsetneq\{i\}_{i=1}^{n}$ for any $I$, and
+Let $K$ be an integer between $1$ and $n$. A $\mathbf{K}$-clustering of $\{i\}_{i=1}^{n}$ is a set $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ such that $C_{I}\subseteq\{1,\ldots,n\}$ for all $I$, and
   1. (a)
@@
     $C_{I}\neq\varnothing\textnormal{ for all }I\in\{1,\ldots,K\}$.
-The total number of clusterings of an $n$-element set grows according to the Bell numbers (Fig. 3), illustrating the combinatorial complexity of the clustering problem.
-We will now define the silhouette width, which was introduced in 1987 by P.J. Rousseeuw [ogsilhouette]. We employ similar notation.
+Remark. The case $K=1$ corresponds to the trivial clustering $\mathcal{C}_{1}=\{\{1,\ldots,n\}\}$, while $K=n$ corresponds to the trivial partition into singletons.
+In the continuation of the article, we only consider $K>1$ for convenience.
+Although we emphasize that the method proposed in Section 4 does not require such an assumption, and can be applied to datasets with a single cluster as well.
+The total number of clusterings of an $n$-element set grows according to the Bell numbers $B_{n}$. For instance, the number of possible partitions of only 26 objects is $B_{26}=49631246523618756274$.
+We will now define the silhouette width, which was introduced in 1987 by P.J. Rousseeuw [18]. We employ similar notation.
 Definition 3 (Silhouette).
-Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{1,\dots,n\}$ with dissimilarity matrix $\Delta$. The silhouette width for data point $i\in C_{I}$ is 0 if $|C_{I}|=1$; otherwise it is
-$s(i|\mathcal{C}_{K},\Delta)\coloneqq\frac{b(i)-a(i)}{\max\{a(i),b(i)\}},$
+Let $\mathcal{C}_{K}$ be a $K$-clustering of the dataset $\{1,\dots,n\}$ with dissimilarity matrix $\Delta$. The silhouette width for data point $i\in C_{I}$ is defined as
+$s(i\mid\mathcal{C}_{K},\Delta)\coloneqq\begin{cases}\dfrac{b(i)-a(i)}{\max\{a(i),b(i)\}}&\text{ if }|C_{I}|>1\land\max\{a(i),b(i)\}>0,\\[3.00003pt] 0&\text{ if }\left(|C_{I}|>1\land\max\{a(i),b(i)\}=0\right)\lor|C_{I}|=1\end{cases}$
 where
 $a(i)\coloneqq\frac{1}{|C_{I}|-1}\sum_{j\in C_{I}\setminus\{i\}}\Delta_{ij},\;\;\;\text{and}\;\;\;b(i)\coloneqq\min_{J\neq I}\left(\frac{1}{|C_{J}|}\sum_{j\in C_{J}}\Delta_{ij}\right).$
-The average silhouette width (ASW) is defined as
-$ASW(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{n}\sum_{i=1}^{n}s(i|\mathcal{C}_{K},\Delta).$
+The average silhouette width ($\operatorname{ASW}$) is defined as
+$\operatorname{ASW}(\mathcal{C}_{K},\Delta)\coloneqq\frac{1}{n}\sum_{i=1}^{n}s(i\mid\mathcal{C}_{K},\Delta).$
 Note that the silhouette width of $i$ may be written as
@@
 Consequently, $1\leq s(i\,|\,\mathcal{C}_{K},\Delta)\leq 1$.
-Individual silhouette widths provide a valuable graphical diagnostic, but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [medoid].
+Individual silhouette widths provide a useful graphical diagnostic, but for large $n$ it is customary to summarize the clustering quality by $\operatorname{ASW}$ [14].
 A common heuristic is to run a clustering algorithm for several choices of $K$ and select the value that maximizes $\operatorname{ASW}$.
 When even that maximum is low, it is possible that the limitation stems from the intrinsic structure of the dissimilarity data.
-This raises the natural question: given a dissimilarity matrix, what is the highest possible $\mathrm{ASW}$?
-To discuss this question, let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering.
-Definition 4 (Silhouette-optimal clustering).
-Let $\Delta$ be the dissimilarity matrix of the dataset $\{1,\ldots,n\}$. We define
-$\mathfrak{C}\coloneqq\{\mathcal{C}_{K}:\mathcal{C}_{K}\text{ is a $K$-clustering of $\{1,\ldots,n\}$}\}$
-to be the set of all possible clusterings. Any
-$\mathcal{C}^{*}\in\operatorname*{argmax}_{\mathcal{C}_{K}\in\mathfrak{C}}\{\mathrm{ASW}(\mathcal{C}_{K},\Delta)\}$
-is called a silhouette-optimal clustering.
+This raises the natural question: given a dissimilarity matrix, what is the highest possible $\operatorname{ASW}$?
+In the remainder of the paper, we let $\mathcal{C}^{*}$ denote a silhouette-optimal clustering.
+That is, if $\mathcal{C}_{K}$ is a $K$-clustering of $\Delta$ then $\mathrm{ASW}(\mathcal{C}_{K},\Delta)\leq\mathrm{ASW}(\mathcal{C}^{*},\Delta)$.
 4 A data-dependent upper bound of the silhouette
-We propose a data-dependent ceiling on the ASW that can be computed before any clustering is performed.
-For each data point, we derive a sharp upper bound of its silhouette width.
-Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and there exists a clustering in which the bound is exactly achieved.
-Averaging these per-point ceilings produces a single, dataset-level upper bound on the ASW.
+We propose a data-dependent ceiling on the $\operatorname{ASW}$ that can be computed before any clustering is performed.
+For each data point, we derive an upper bound of its silhouette width.
+Averaging these per-point ceilings produces a single, dataset-level upper bound of the $\operatorname{ASW}$.
 The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and the proposed algorithm supports optional minimum cluster size constraints to further tighten the bound.
@@
 $s(i|\mathcal{C}_{K},\Delta)$.
+For our purposes, it will be convenient to work with the off-diagonal entries of $\Delta$, sorted row-wise in ascending order.
+We denote this transformed dissimilarity matrix by $\hat{\Delta}$ ($\hat{\Delta}\in\mathbb{R}^{n\times n-1}$).
+In other words, we have for every $i$ $\{\Delta_{ij}\}_{j=1}^{n}\setminus\{\Delta_{ii}\}=\{\hat{\Delta}_{ij}\}_{j=1}^{n-1}$ and $\hat{\Delta}_{i1}\leq\hat{\Delta}_{i2}\leq\cdots\leq\hat{\Delta}_{i(n-1)}$.
+Constructing $\hat{\Delta}$ requires sorting each of the $n$ rows of $\Delta$, which costs $\mathcal{O}(n^{2}\log n)$ time in total.
 For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$.
@@
 These two observations are combined to construct a value that is guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$.
-We formalize this reasoning through the $\Lambda$-quotient of point $i$.
+We formalize this reasoning through the $k$-quotient of point $i$.
 Definition 4.
-Let $\Lambda\in\mathbb{N}$ with $1\leq\Lambda\leq n-1$. The $\Lambda$-quotient of data point $i$ is
-$q(i,\Delta,\Lambda)\coloneqq\begin{cases}\dfrac{n-\Lambda}{\Lambda-1}\cdot\dfrac{\sum_{j=1}^{\Lambda-1}\hat{\Delta}_{ij}}{\sum_{j=\Lambda}^{n-1}\hat{\Delta}_{ij}}&\textnormal{if }\Lambda>1,\\[4.30554pt] 1,&\textnormal{if }\Lambda=1.\end{cases}$
+Let $k\in\mathbb{N}$ with $1\leq k\leq n-1$. The $k$-quotient of data point $i$ is
+$q(i,\Delta,k)\coloneqq\begin{cases}\dfrac{n-k}{k-1}\cdot\dfrac{\sum_{j=1}^{k-1}\hat{\Delta}_{ij}}{\sum_{j=k}^{n-1}\hat{\Delta}_{ij}}&\textnormal{if }k>1,\\[4.30554pt] 1,&\textnormal{if }k=1.\end{cases}$
 Recall the following basic fact.
@@
 Furthermore, let
-$f(i,\Delta)\coloneqq\min_{1\leq\Lambda\leq n-1}\{q(i,\Delta,\Lambda)\}$
-denote a minimal $\Lambda$-quotient. Then
+$f(i,\Delta)\coloneqq\min_{1\leq k\leq n-1}\{q(i,\Delta,k)\}$
+denote a minimal $k$-quotient. Then
 $s(i|\mathcal{C}_{K},\Delta)\leq 1-f(i,\Delta).$
@@
 Taking the mean of all silhouette width upper bounds from Proposition 1 results in the following inequality, which is the main observation of this paper:
-$ASW(\mathcal{C}^{*},\Delta)\leq 1-\frac{1}{n}\sum_{i}f(i,\Delta).$ (1)
-We remark that the right-hand side in 1 depends only on the underlying data $\Delta$.
-Crucially, we note that $1-\frac{1}{n}\sum_{i}f(i,\Delta)\leq 1$.
-Moreover, since any $\Lambda$-quotient naturally corresponds to a $2$-clustering of the dataset, the bound in Proposition 1 is indeed sharp.
-Finally, imposing the narrower definition
-$f_{\kappa}(i,\Delta)\coloneqq\min_{\kappa\leq\Lambda\leq n-\kappa}\{q(i,\Delta,\Lambda)\}$
-for some $\kappa\in\{1,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$.
-Often, as we shall see, this leads to strictly tighter limits.
+$\operatorname{ASW}(\mathcal{C}^{*},\Delta)\leq 1-\frac{1}{n}\sum_{i}f(i,\Delta).$ (1)
+We remark that the right-hand side in 1 depends only on the underlying data $\Delta$. Crucially, we note that $1-\frac{1}{n}\sum_{i}f(i,\Delta)\leq 1$.
+We stress that the right-hand side in 1 is not, in general, a sharp upper bound of the ASW.
+However, for a single data point $i$, the quantity $1-f(i,\Delta)$ is a sharp upper bound of its individual silhouette width in the sense that the individual silhouette width of the data point can never exceed the bound, and there exists a clustering in which the bound is achieved exactly.
+Indeed, considering the $2$-clustering where $i$ is grouped with its k1k-1 nearest neighbors and all remaining points are grouped into a single separate cluster yields $s(i|C_{K},\Delta)=1-f(i,\Delta)$.
+Imposing the narrower definition
+$f_{m}(i,\Delta)\coloneqq\min_{m\leq k\leq n-m}\{q(i,\Delta,k)\}$
+for some $m\in\{1,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq m$.
+By letting $\mathcal{C}^{*}_{m}$ denote a silhouette-optimal clustering within this solution space, we have $\operatorname{ASW}(\mathcal{C}^{*}_{m},\Delta)\leq 1-\frac{1}{n}\sum_{i}f_{m}(i,\Delta)$.
+In real-world applications, analysts sometimes need to enforce minimum cluster size constraints to avoid too small clusters [24, Höppner2008].
+From this perspective, the quantity $1-\frac{1}{n}\sum_{i}f_{m}(i,\Delta)$ does not aim to approximate the unconstrained global upper bound in 1, but rather provides an upper bound tailored to the family of clusterings that a practitioner is willing to consider.
+This allows for a meaningful benchmark in settings where small clusters are inadmissible, and offers flexibility in situations where minimum cluster sizes are part of the problem specification.
 4.1 Algorithm
@@
 1
-Input: $\kappa$;
-// Smallest allowed cluster size, $1\leq\kappa\leq\lfloor n/2\rfloor$
+Input: $m$;
+// Smallest allowed cluster size, $1\leq m\leq\lfloor n/2\rfloor$
 2
-Output: $A\in\mathbb{R}^{n}$;
-// Array of individual upper bounds
+Output: $S\in\mathbb{R}^{n}$;
+// Vector of point-wise upper bounds
 3
-4 Initialize array $A\leftarrow[\;]$ ;
+4 Initialize $S\leftarrow[\;]$ ;
 5 Compute $\hat{\Delta}$ from $\Delta$;
 6 for $i=1,\ldots,n$ do
-7 $y\leftarrow\sum_{j=\kappa}^{n-1}\hat{\Delta}_{i,j}$ ;
+7 $y\leftarrow\sum_{j=m}^{n-1}\hat{\Delta}_{i,j}$ ;
 8 $x\leftarrow 0$ ;
  $q\leftarrow 1$ ;
- // Initialize $\Lambda$-quotient
-9 if $\kappa>1$ then
-10    $x\leftarrow\sum_{j=1}^{\kappa-1}\hat{\Delta}_{i,j}$ ;
-11    $q\leftarrow\frac{x}{\kappa-1}\big/\frac{y}{n-\kappa}$;
+ // Initialize $k$-quotient
+9 if $m>1$ then
+10    $x\leftarrow\sum_{j=1}^{m-1}\hat{\Delta}_{i,j}$;
+11    $q\leftarrow\frac{x}{m-1}\big/\frac{y}{n-m}$;
 12
-13 for $\Lambda=\kappa+1,\ldots,n-\kappa$ do
-14    $x\leftarrow x+\hat{\Delta}_{i,\Lambda-1}$;
-15    $y\leftarrow y-\hat{\Delta}_{i,\Lambda-1}$;
-16    $q_{\text{candidate}}\leftarrow\dfrac{x}{\Lambda-1}\big/\dfrac{y}{n-\Lambda}$;
+13 for $k=m+1,\ldots,n-m$ do
+14    $x\leftarrow x+\hat{\Delta}_{i,k-1}$;
+15    $y\leftarrow y-\hat{\Delta}_{i,k-1}$;
+16    $q_{\text{candidate}}\leftarrow\dfrac{x}{k-1}\big/\dfrac{y}{n-k}$;
 17    if $q_{\text{candidate}}<q$ then
        $q\leftarrow q_{\text{cand}}$ ;
-       // Update minimal $\Lambda$-quotient
+       // Update minimal $k$-quotient
 18
 19
-20  Append 1q1-q to AA;
+20  Append 1q1-q to $S$;
 21
-return AA
+return $S$
 Algorithm 1 Silhouette upper bound (pointwise)
-The workflow for computing the proposed upper bound is illustrated in Fig.4.
-Following this workflow, Algo.
-1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the solution space $\min_{I}|C_{I}|\geq\kappa$.
-The array $A$ collects these values across all points, and its average equals the dataset-level upper bound.
-Time complexity
-The formation of $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$.
+Algorithm 1 computes for each data point $i$ the quantity $1-f_{m}(i,\Delta)$, which are sharp upper bounds of the silhouette widths in the solution space $\min_{I}|C_{I}|\geq m$.
+The vector $S$ collects these values across all points, and its average equals the dataset-level upper bound.
+As we shall see in the Experiments section, the data point-level upper bounds provide a nice complement to standard silhouette visualizations, where the achieved point-wise silhouette widths are typically displayed next to each other in a bar graph.
+Scalability
+The formation of $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$ time.
 The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall runtime is $\mathcal{O}(n^{2}\log n)$.
-This is slightly slower than computing the $\mathrm{ASW}$, which takes $\mathcal{O}(n^{2})$ time [medoid].
+This is slightly slower than computing the $\operatorname{ASW}$, which takes $\mathcal{O}(n^{2})$ time [14].
+Since the algorithm operates on the entire $n\times n$ distance matrix, the space complexity of Algorithm 1 is $\mathcal{O}(n^{2})$.
+This quadratic memory requirement quickly becomes the dominant bottleneck.
+For example, a dataset with $10^{5}$ samples produces $10^{10}$ pairwise distances; storing these in float32 requires 40 GB of memory.
+In practice, this limits the algorithm to datasets with at most a few tens of thousands of points on standard hardware.
+There are ways to mitigate this in large-scale settings.
+For example, one may approximate the upper bound by computing it on a random subset of the samples.
+Another possibility is to compute and sort distances on the fly rather than materializing the entire matrix.
+A more thorough investigation of such scalable variants is an interesting direction for future work, but it lies beyond the scope of this paper.
 Output summaries
-Given $\Delta$ and $\kappa=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $A$ from Algo.
-1.
+Given $\Delta$ and $m=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $S$ from Algorithm 1.
 Then $\texttt{UB}(\Delta)$ corresponds to the right-hand side of Eq.
@@
 We refer to this as the global upper bound.
-Choosing $\kappa>1$ imposes the restriction that no cluster is allowed to contain fewer than $\kappa$ objects.
-Let $\texttt{UB}_{\kappa}(\Delta)$ denote the $\mathrm{ASW}$ upper bound for this constrained solution space.
-We will refer to $\texttt{UB}_{\kappa}(\Delta)$ as an adjusted upper bound.
-Remark
-The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$. Its tightness is data-dependent.
-Use cases
-In theory, when sharp, the upper bound $\texttt{UB}(\Delta)$ can be used to confirm that an empirical ASW has reached the global optimum.
-Even when loose, it increases the interpretability of empirical measurements.
-In real-world scenarios where the global upper bound appears loose, we speculate that adjusted versions can provide more practical use.
-The individual silhouette width upper bounds generated by Algo.
-1 provide a neat complement to standard silhouette visualizations, where the point-wise silhouette widths are typically displayed next to each other in a bar graph.
-In the Experiments section, we examine these use cases.
-Figure 4: Workflow for computing the proposed $\mathrm{ASW}$ upper bound.
+Choosing $m>1$ imposes the restriction that no cluster is allowed to contain fewer than $m$ objects.
+Let $\texttt{UB}_{m}(\Delta)$ denote the $\operatorname{ASW}$ upper bound for this constrained solution space.
+We will refer to $\texttt{UB}_{m}(\Delta)$ as a constrained upper bound.
 4.2 Extension to the macro-averaged silhouette
-The standard $\mathrm{ASW}$ can produce misleading results when the clusters vary greatly in size.
-In [john], John Pavlopoulos, Georgios Vardakas, and Aristidis Likas show that the so-called macro-averaged silhouette is more robust to cluster-size imbalance.
+The standard $\operatorname{ASW}$ can produce misleading results when the clusters vary greatly in size.
+John Pavlopoulos, Georgios Vardakas, and Aristidis Likas [16] show that the so-called macro-averaged silhouette is more robust to cluster size imbalance.
 For clustering $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$, the macro-averaged silhouette is defined as
@@
 Thus, an efficiently computable data-dependent upper bound is of interest as a relevant benchmark.
-Let $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ be any clustering from our constrained solution space, and let $\texttt{ub}_{i}$ denote the silhouette width upper bound of data point $i$, i.e.
-$s(i|\mathcal{C}_{K},\Delta)\leq\texttt{ub}_{i}$.
+Let $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ be any clustering from our constrained solution space, and let $\texttt{ub}_{i}$ denote the silhouette width upper bound of data point $i$, i.e., $s(i|\mathcal{C}_{K},\Delta)\leq\texttt{ub}_{i}$.
 Moreover, let $\{\texttt{ub}_{\sigma(i)}\}_{i=1}^{n}$ denote the upper bounds sorted in ascending order, and let $\{N_{I}\}_{I=1}^{K}$ denote the assumed cluster sizes in descending order.
@@
 The last expression is independent of any particular cluster assignment and therefore provides a valid upper bound on the macro-averaged silhouette across all clusterings within the constrained solution space.
-This formulation is particularly useful in applications where cluster sizes are known or can be reasonably estimated.
+This formulation is particularly useful in applications where cluster sizes are known or can be reasonably estimated, for example in balanced clustering [15].
 Even in the absence of such assumptions, the resulting bound serves as a sharper and more informative benchmark for evaluating empirical clustering results.
-5 Experiments, results & discussion
-Table 1: Real-world datasets. Datasets used in the experiments and their characteristics. The Metric column indicates the distance metrics used.
+Because the macro-averaged silhouette weights clusters equally, it carries the danger of favoring solutions with a few small outlier clusters while masking poor structure in the remaining parts.
+It should therefore be used only when equal importance of all clusters is appropriate.
+Moreover, we do not recommend routine comparisons with ASW via heuristics such as $\max(\mathrm{ASW},S_{macro})$.
+Informed choices should instead be made on a case-by-case basis.
+We stress that this extension is not a central focus of the paper. Our aim here is only to illustrate an idea, rather than to conduct a full empirical study of this variant.
+5 Experiments
+We evaluate the $\operatorname{ASW}$ upper bound by applying it to a variety of datasets, both synthetic and real.
+Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity.
+The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound which contains all preprocessing steps, dissimilarity constructions, and scripts required to reproduce the reported results.
+For a given dissimilarity matrix, an obvious desideratum of the proposed upper bound is that it is relatively close to the maximum $\operatorname{ASW}$.
+However, the maximum $\operatorname{ASW}$ is of course typically unknown.
+This presents a fundamental challenge in the evaluation process.
+To remedy this, we must try to establish a nearly tight lower bound of $\operatorname{ASW}(\mathcal{C}^{*},\Delta)$.
+To this end, we use PAMSIL clustering by default to obtain an empirical $\operatorname{ASW}$.
+The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies.
+PAMSIL is a suitable algorithm since its objective is to maximize $\operatorname{ASW}$.
+When dealing with larger datasets, we use FastMSC and DynMSC instead.
+Given an empirical $\operatorname{ASW}$ value, say $\widehat{S}>0$, the performance metric of interest is the relative error, defined as
+$\frac{\operatorname{ASW}(\mathcal{C}^{*},\Delta)-\widehat{S}}{\operatorname{ASW}(\mathcal{C}^{*},\Delta)}=1-\frac{\widehat{S}}{\operatorname{ASW}(\mathcal{C}^{*},\Delta)},$
+where $\mathcal{C}^{*}$ denotes a clustering that achieves the global maximum.
+Since $\operatorname{ASW}(\mathcal{C}^{*},\Delta)$ is unknown in practice, we use the global upper bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
+$1-\frac{\widehat{S}}{\operatorname{ASW}(\mathcal{C}^{*},\Delta)}~\leq~1-\frac{\widehat{S}}{\texttt{UB}(\Delta)}.$
+We refer to this quantity as the worst-case relative error.
+Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical $\operatorname{ASW}$ is guaranteed to be within $\epsilon$ of the true maximum.
+Instead of employing this measure to compare clustering algorithms, we will use it to evaluate the tightness of the upper bound relative to the $\operatorname{ASW}$ obtained by the chosen algorithm.
+The results of both the global upper bound and the constrained upper bound are reported.
+In our experiments, we do not impose external constraints on the minimum cluster size.
+Instead, for each empirical clustering solution (e.g., produced by PAMSIL), we simply record its smallest cluster size and compute $\texttt{UB}_{m}({\Delta})$ using this value of $m$.
+This approach does not enforce a constraint but allows us to observe how the constrained upper bound behaves under the cluster size patterns that arise in practice.
+5.1 Synthetic data
+In this section, we look at the proposed upper bound applied to synthetic data generated by the make_blobs function from the scikit-learn library [17].
+In this idealized setting, we also demonstrate why a complete scan of the $k$-values is necessary when constructing $f(i,\Delta)$ in Proposition 1.
+Indeed, it is reasonable to expect that $k=2$ often minimizes $q(i,\Delta,k)$.
+Lastly, we use synthetic dissimilarity matrices to evaluate the runtime of Algorithm 1.
+In the following sections, we evaluate the usefulness of the upper bound in more realistic settings.
+For dataset notation, we follow [8].
+Each dataset is labeled as n_samples-n_features-centers-cluster_std, indicating the parameter values used to generate the points.
+For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion.
+For these datasets, we use Euclidean distance.
+Table 2: Experimental results on synthetic data. Comparison of clustering performance and the proposed upper bound. Avg. distance indicates the average pairwise distance between cluster centers.
+Dataset $K$ ARI AMI $\operatorname{ASW}$ $\texttt{UB}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}(\Delta)}$ Avg. distance
+400-64-5-6 5 1.0 1.0 $0.249$ $0.376$ $0.339$ $65.2$
+400-64-2-2 2 1.0 1.0 $0.673$ $0.673$ $0.000$ $64.5$
+400-128-7-3 7 1.0 1.0 $0.522$ $0.566$ $0.078$ $92.3$
+1000-300-5-2 5 1.0 1.0 $0.668$ $0.678$ $0.015$ $143.7$
+10000-32-20-2 20 1.0 1.0 $0.626$ $0.774$ $0.192$ $47.0$
+10000-1024-20-4 20 1.0 1.0 $0.417$ $0.454$ $0.081$ $263.1$
+In Table 2, we apply the upper bound to six synthetic datasets and compare the results with PAMSIL clustering.
+For dataset 400-64-2-2, the upper bound at $0.673$ confirms that PAMSIL reaches a globally optimal silhouette score.
+For three of the remaining datasets, $\texttt{UB}(\Delta)$ proves that the PAMSIL $\operatorname{ASW}$ is within 8 $\%$ of the global maximum.
+For dataset 400-64-5-6 with $\texttt{UB}(\Delta)=0.376$, we record an empirical $\operatorname{ASW}$ of $0.249$.
+In this case, it remains unclear whether further substantial improvements of the $\operatorname{ASW}$ are possible.
+However, interpreting the attained value in the interval [1,0.376][-1,0.376] enables more well-founded conclusions compared to using the standard range [1,1][-1,1].
+Lastly, the reported ARI and AMI scores tell us that PAMSIL finds the ground truth clustering in all six datasets.
+Figure 3 illustrates why it is necessary to scan all values of $k$ when constructing the upper bound in Proposition 1.
+In this example, we have picked a random data point $i$ from the dataset 1000-300-5-2, and the graph shows the quantity $1-q(i,\Delta,k)$ over a range of $k$-values, demonstrating that, for example, $k=2$ does not correspond to a valid bound in this case.
+Instead, the minimum of $q(i,\Delta,k)$ occurs at $k=200$, which coincides with the generated cluster sizes.
+Although in real-world datasets with less clearly separated clusters, we generally expect the difference between the average distance to the n2n-2 farthest neighbors and the nearest-neighbor distance to be sufficiently large that $k=2$ minimizes $q$.
+We suspect that this behavior often holds true even for data points whose cluster size in $\mathcal{C}^{*}$ is much larger than 2.
+Assuming that this speculation is true, it would be very unlikely that $\texttt{UB}(\Delta)$ is actually close to sharp in real-world settings.
+Figure 3: For a random data point in the synthetic dataset 1000-300-5-2, we show the quantity $1-q(\Delta,k)$ for a range of $k$-values.
+The red dashed horizontal line shows the point’s silhouette width achieved by PAMSIL clustering with $K=5$.
+The maximum of $1-q(\Delta,k)$ is the upper bound from Proposition 1.
+The maximum is given at $k=200$, which corresponds to the cluster sizes in this dataset.
+Figure 4 shows runtime measurements for Algorithm 1 (mm=1).
+Additionally, we include the runtime of constructing $\hat{\Delta}$ from $\Delta$, i.e., the second row in Algorithm 1.
+As a reference, we also report the time needed to compute the dissimilarity matrix $\Delta$ (although $\Delta$ is assumed to be precomputed in Algorithm 1).
+The experiments were conducted on an Apple MacBook Pro (M1, 2020) with 16GB of RAM, and the software environment was Python 3.10.
+Distance matrices of varying size and float32 data type were generated through the NumPy and scikit-learn packages.
+Specifically, we used the random.randn function to generate $n$ samples with $32$ features (seed fixed to 0), and the metrics.pairwise_\_distances function to generate all pairwise distances.
+The runtime of Algorithm 1 does not depend on the degree of clustering structure present in the dataset, which is why make_\_blobs was not used for this task.
+For each distance matrix, the programs were executed five times.
+The graph shows the average runtimes for each number of samples $n$, and the error bars indicate one standard deviation.
+It is clear from the figure that in our implementation, a substantial amount of the runtime of Algorithm 1 is spent on sorting each of the $n$ rows of $\Delta$ to obtain $\hat{\Delta}$.
+Figure 4: Runtime experiments on synthetic data (log-linear plot).
+5.2 UCI datasets
+We evaluate our method on eight real-world datasets from the UCI Machine Learning Repository [13], commonly used in clustering and classification research.
+Table 3 summarizes their key characteristics along with the distance metrics employed.
+Standard preprocessing steps, including feature scaling and outlier removal, are applied where appropriate.
+Each dataset provides a known number of classes, which we use to determine the number of clusters $K$.
+When all class labels are available, we report the ARI and AMI to verify that these datasets represent meaningful clustering structures.
+Table 3: UCI datasets. Datasets used in the experiments and their characteristics. The Metric column indicates the distance metrics used.
 Dataset Title Instances Features Classes Field Metric
@@
 Wine Wine 178 13 3 Chemistry Manhattan
-Table 2: Experimental results on synthetic data. Comparison of clustering performance and the proposed upper bound.
-Dataset $K$ ASW $\texttt{UB}(\Delta)$ Worst-case rel. error
-400-64-5-6 5 $0.249$ $0.376$ $0.339$
-400-64-2-2 2 $0.673$ $0.673$ $0.000$
-400-128-7-3 7 $0.522$ $0.566$ $0.078$
-1000-300-5-2 5 $0.668$ $0.678$ $0.015$
-10000-32-20-2 20 $0.626$ $0.774$ $0.192$
-10000-1024-20-4 20 $0.417$ $0.454$ $0.081$
-We evaluate the ASW upper bound by applying it to a variety of datasets, both synthetic and real-world.
-Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity.
-The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound.
-An essential part of the evaluation is to establish a nearly tight lower bound for $\mathrm{ASW}(C^{*},\Delta)$.
-To this end, we use PAMSIL clustering to obtain an empirical $\mathrm{ASW}$.
-The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies.
-PAMSIL is a suitable algorithm since its objective is to maximize ASW.
-Given an empirical ASW value, say $\widehat{S}$, the performance metric of interest is the relative error, defined as
-$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)},$
-where $C^{*}$ denotes a clustering that achieves the global maximum.
-Since $\mathrm{ASW}(C^{*},\Delta)$ is unknown in practice, we use the global upper bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
-$\frac{\mathrm{ASW}(C^{*},\Delta)-\widehat{S}}{\mathrm{ASW}(C^{*},\Delta)}~\leq~\frac{\texttt{UB}(\Delta)-\widehat{S}}{\texttt{UB}(\Delta)}.$
-We refer to this quantity as the worst-case relative error.
-Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical ASW is guaranteed to be within $\epsilon$ of the true maximum.
-Synthetic datasets
-We follow [gaussian] in notation and approach.
-Synthetic datasets are generated using the make_blobs function from scikit-learn [scikit].
-Each dataset is labeled as n_samplesn_featurescenterscluster_stdn\_samples-n\_features-centers-cluster\_std, indicating the parameter values used to generate the points.
-For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion.
-Finally, for these datasets, we use Euclidean distance.
-Real-world datasets
-We evaluate our method on eight real-world datasets from the UCI Machine Learning Repository [ucidata], commonly used in clustering and classification research.
-Table 1 summarizes their key characteristics along with the distance metrics employed.
-Standard preprocessing steps, including feature scaling and outlier removal, are applied where appropriate.
-Each dataset provides a known number of classes, which we use to determine the number of clusters $K$.
-When all class labels are available, we report the ARI and AMI to verify that these datasets represent meaningful clustering structures.
-In that way, we ensure that our experiments are conducted in natural clustering contexts.
-Table 3: Experimental results on real data. Comparison of clustering performance and the proposed upper bound.
-Performance of PAMSIL clustering
-Comparison with
-global upper bound
-Comparison with
-adjusted upper bound
-Dataset $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI ASW $\texttt{UB}(\Delta)$ Worst-case rel. error $\texttt{UB}_{\kappa}(\Delta)$ Worst-case rel. error
+Table 4: Experimental results on UCI datasets. Comparison of clustering performance and the proposed upper bound.
+Performance of PAMSIL clustering Comparison with global upper bound Comparison with constrained upper bound
+Dataset $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI $\operatorname{ASW}$ $\texttt{UB}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}(\Delta)}$ $\texttt{UB}_{m}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}_{m}(\Delta)}$
 Ceramic 2 44 $1.000$ $1.000$ $0.358$ $0.621$ $0.423$ $0.403$ $0.112$
@@
 Wine 3 53 $0.884$ $0.864$ $0.315$ $0.649$ $0.514$ $0.450$ $0.298$
-Figure 5: Silhouette plots for the eight real-world datasets.
-PAMSIL clustering with $K$ equal to the ground truth number of classes; ASW compared against the global upper $\texttt{UB}(\Delta)$ and the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size.
-The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds.
+Table 4 examines the performance of PAMSIL on the eight real-world datasets and compares it with the upper bounds.
+In these cases, the gap between the empirical $\operatorname{ASW}$ and the global upper bound is generally quite large.
+It is unclear whether the reason for this is that the attained $\operatorname{ASW}$ values are far from $\operatorname{ASW}(\mathcal{C}^{*},\Delta)$, or that the upper bounds are rather loose, or both.
+However, the constrained upper bounds $\texttt{UB}_{m}(\Delta)$ are considerably tighter for most datasets.
+Here, $m$ is set to the minimum cluster size in the PAMSIL solution.
+For datasets Ceramic, Customers, Rna, Wdbc, and Wine, $\texttt{UB}_{m}(\Delta)$ proves that the $\operatorname{ASW}$-values attained by PAMSIL are within 30 $\%$ of the corresponding optimums (within the solution spaces constrained by $m$).
+Figure 5: Silhouette plots for the UCI datasets.
+PAMSIL clustering with $K$ equal to the ground truth number of classes; $\operatorname{ASW}$ compared against the global upper $\texttt{UB}(\Delta)$ and the constrained upper bound $\texttt{UB}_{m}(\Delta)$, with $m$ being the smallest cluster size.
+The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding constrained upper bounds, while the transparent areas indicate the corresponding global upper bounds.
 Figure 6: PAMSIL performance for varying $K$ and comparison with upper bounds.
-Results
-In Table 2, we apply the upper bound to six synthetic datasets and compare the results with PAMSIL clustering.
-For dataset 400-64-2-2, the upper bound at $0.673$ confirms that PAMSIL reaches a globally optimal silhouette score.
-For three of the remaining datasets, $\texttt{UB}(\Delta)$ proves that the PAMSIL ASW is within 8 $\%$ of the global maximum, that is, essentially optimal.
-For dataset 400-64-5-6 with $\texttt{UB}(\Delta)=0.376$, we record an empirical ASW of $0.249$.
-In this case, it remains unclear whether further substantial improvements of the ASW are possible.
-However, interpreting the attained value in the interval [1,0.376][-1,0.376] enables more well-founded conclusions compared to using the standard range [1,1][-1,1].
-Table 3 examines the performance of PAMSIL on the eight real-world datasets and compares it with the upper bounds.
-In these cases, the global upper bounds provide information that is less actionable.
-It is unclear whether the reason for this is that the attained ASW values are far from $\mathrm{ASW}(C^{*},\Delta)$, or that the upper bounds are somewhat loose, or both.
-However, the adjusted upper bounds $\texttt{UB}_{\kappa}(\Delta)$ are considerably tighter for most datasets.
-Here, $\kappa$ is set to the minimum cluster size in the PAMSIL solution.
-For datasets Ceramic, Customers, Rna, Wdbc and Wine, $\texttt{UB}_{\kappa}(\Delta)$ proves that the ASW-values attained by PAMSIL are within 30 $\%$ of the corresponding optimums.
-Fig.
-5 shows classical silhouette visualizations corresponding to the PAMSIL clusters, complemented with upper bounds of the individual silhouette widths.
-The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds.
-For the adjusted upper bounds, $\kappa$ is set to the smallest cluster size in the generated solution.
+As for the global version, the lowest upper bound we recorded was for the Rna dataset, for which PAMSIL reached an ASW of $0.225$ and the upper bound was $0.419$.
+In contrast, the highest upper bound we recorded was for the Customers dataset, for which PAMSIL reached an ASW of $0.4$ and the upper bound was $0.951$.
+The smallest and largest worst-case relative errors were for the Ceramic dataset ($0.423$) and the Heart Statlog dataset ($0.731$), respectively.
+We observe that the gaps between $\texttt{UB}(\Delta)$ and $\texttt{UB}_{m}(\Delta)$ are rather large for most datasets.
+For instance, the Heart Statlog dataset showed $\texttt{UB}(\Delta)=0.6$ and $\texttt{UB}_{83}(\Delta)=0.28$ — a 53 $\%$ decrease.
+Earlier in the text, when discussing Figure 3, we speculated that, in many real-world datasets, the $k$-quotient will frequently reach its minimum at $k=2$ regardless of the cluster structure in $\mathcal{C}^{*}$.
+If $q(\Delta,m)\gg q(\Delta,2)$ this naturally leads to large gaps.
+On the one hand, the constrained upper bound is an important feature of the proposed method that can lead to significantly sharper insights than those provided by the global upper bound.
+On the other hand, implementing an unjustified constraint by mistake can cause incorrect conclusions.
+Therefore, we recommend careful consideration when choosing the parameter $m$.
+Figure  5 shows classical silhouette visualizations corresponding to the PAMSIL clusters, complemented with upper bounds of the individual silhouette widths.
+The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding constrained upper bounds, while the transparent areas indicate the corresponding global upper bounds.
+For the constrained upper bounds, $m$ is set to the smallest cluster size in the generated solution.
 This method provides a visual comparison of the empirical silhouette and the highest possible value at the data point level.
-In Fig.
-6 we show, for each real-world dataset and for each $K\in\{2,\ldots,15\}$, the attained ASW value compared to the global ASW upper bound.
-Additionally, at each $K$, we compare the ASW to the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$ with $\kappa$ set to the minimum cluster size in the corresponding solution.
+In Figure  6 we show, for each dataset and for each $K\in\{2,\ldots,15\}$, the attained $\operatorname{ASW}$ value compared to the global $\operatorname{ASW}$ upper bound.
+Additionally, at each $K$, we compare the $\operatorname{ASW}$ to the constrained upper bound $\texttt{UB}_{m}(\Delta)$ with $m$ set to the minimum cluster size in the corresponding solution.
 Lastly, we include, at each $K$, the recorded macro-averaged silhouette and compare this value with the upper bound derived in Section 4.2.
-Discussion
-We summarize with some reflections on the practical relevance of the proposed upper bound.
-Replacing the standard $\mathrm{ASW}$ range [1,1][-1,1] with the data-dependent interval $1,\texttt{UB}(\Delta)]$ enhances the interpretability of observed $\mathrm{ASW}$ values.
-However, when the empirical $\mathrm{ASW}$ lies far below the upper bound, it remains unclear whether further improvement is actually possible.
-We suspect that this gap often persists for the global upper bound, even when the empirical $\mathrm{ASW}$ is close to $\mathrm{ASW}(C^{*},\Delta)$.
-In many real-world applications, however, analysts impose minimum cluster-size constraints to avoid overly small or degenerate clusters.
-By incorporating such constraints through $\kappa$, the refined bound $\texttt{UB}_{\kappa}(\Delta)$ better reflects the relevant solution space and provides more meaningful guidance in practice.
-As demonstrated in Table 3, $\texttt{UB}_{\kappa}(\Delta)$ ensures that the PAMSIL solutions are within 30 $\%$ of the optimum for five of the examined datasets, indicating strong empirical performance under realistic conditions.
-In such cases, the bound can be used to conclude that further exploration would, at best, yield only marginally better results.
-6 Conclusion
-In this paper, we proposed a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a complementary tool for analyzing clustering quality.
-While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a tighter data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time.
-This upper bound provides a more meaningful reference value.
+This scan over different $K$ reveals that PAMSIL reaches a higher or equally high $\operatorname{ASW}$ at a value of $K$ that is different than the ground truth number of classes for all datasets but Ceramic and Wdbc.
+5.3 The ALOI dataset
+The purpose of this experiment is to study how the proposed upper bound behaves on larger datasets with a large number of clusters.
+In particular, we investigate whether the bound appears informative when the number of classes is large.
+We base our experiments on data derived from the Amsterdam Library of Object Images (ALOI) [21], a large-scale image collection containing objects photographed under varying viewing angles and illumination conditions.
+Each image is represented using HSV/HSB color histograms with different binning resolutions.
+In this representation, each pixel is transformed into hue–saturation–brightness space and assigned to its nearest histogram bin.
+We consider two such representations, namely $5\!\times\!5\!\times\!5$ and $14\!\times\!6\!\times\!6$ binnings, resulting in feature vectors of up to several hundred dimensions.
+In this setting, we use the Manhattan distance metric.
+The datasets have 1000 object classes with at least 100 instances available for each class.
+To obtain a total number of samples that is manageable, we randomly select $40$ instances per class, yielding datasets with 40,00040{,}000 samples in total.
+In this regime, PAMSIL becomes computationally prohibitive, and we therefore rely on FastMSC and DynMSC, which are significantly more scalable while still directly optimizing silhouette-based objectives.
+For these experiments, we report both clustering runtimes and the runtimes required to compute the global and constrained upper bounds.
+Table 5 summarizes the results.
+With $K=1000$, the gap between the global upper bound and the empirical $\operatorname{ASW}$ is rather substantial for both datasets.
+However, applying DynMSC ($2\leq K\leq 50$) reveals that a much higher $\operatorname{ASW}$ is achievable at $K=2$.
+For the $14\!\times\!6\!\times\!6$ dataset with $m=2$ and $m=3$, the constrained upper bounds show slightly higher runtimes than the global upper bounds, which we attribute to runtime variability in the experimental environment.
+We also conduct some small-scale experiments with 2 and 5 classes from the same underlying data.
+For these datasets, we sample 100 instances from each class to obtain datasets of sizes 200 and 500, respectively.
+For instance, $5\!\times\!5\!\times\!5-2$ and $5\!\times\!5\!\times\!5-5$ are datasets derived from the same image representations, but with different number of classes (2 and 5).
+Since these datasets are so small, we can apply PAMSIL and compare it with FastMSC and DynMSC.
+Table 6 summarizes the results.
+It is difficult to draw strong conclusions from Table 5.
+The large-scale setting introduces substantial uncertainty regarding the most appropriate number of clusters, which complicates a precise assessment of the bounds tightness.
+In contrast, Table 6 provides a more controlled comparison.
+The results suggest that the gap between the maximum $\operatorname{ASW}$ and the proposed upper bound increases as the number of clusters in the desirable solution grows, even when moving from just $2$ to $5$ clusters.
+This observation highlights an important limitation of the method: the bound appears to be most informative when the underlying clustering structure involves a relatively small number of clusters.
+Consequently, the proposed upper bound should be interpreted with great caution in settings where the true number of clusters is large, and its practical usefulness is expected to be greatest when a small number of clusters is assumed.
+Table 5: Experimental results on ALOI datasets with 1000 classes.
+Comparison of clustering performance and the proposed upper bound.
+DynMSC scans $2\leq K\leq 50$ and FastMSC scans $K=100,200,\ldots,900$.
+Runtime (RT) is expressed in 1000 seconds.
+Performance of FastMSC and DynMSC clustering Comparison with global upper bound Comparison with constrained upper bound
+Dataset Algorithm $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI $\operatorname{ASW}$ RT $\texttt{UB}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}(\Delta)}$ RT $\texttt{UB}_{m}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}_{m}(\Delta)}$ RT
+5x5x5 FastMSC 1000 2 $0.267$ $0.557$ $0.133$ $1.1$ $0.914$ $0.854$ $0.3$ $0.914$ $0.854$ $0.3$
+FastMSC 100 7 $0.006$ $0.368$ $0.181$ $9.7$ $0.914$ $0.802$ $0.3$ $0.866$ $0.791$ $0.3$
+DynMSC 2 976 $0.000$ $0.009$ $0.502$ $1.6$ $0.914$ $0.451$ $0.3$ $0.618$ $0.188$ $0.3$
+14x6x6 FastMSC 1000 2 $0.292$ $0.589$ $0.142$ $1.1$ $0.901$ $0.842$ $0.2$ $0.901$ $0.842$ $0.3$
+FastMSC 100 3 $0.005$ $0.352$ $0.161$ $9.9$ $0.901$ $0.822$ $0.2$ $0.883$ $0.818$ $0.3$
+DynMSC 2 38 $0.000$ $0.001$ $0.470$ $1.7$ $0.901$ $0.478$ $0.2$ $0.752$ $0.375$ $0.2$
+Table 6: Experimental results on ALOI datasets with a small number of classes. Comparison of clustering performance and the proposed upper bound. DynMSC scans $2\leq K\leq 10$.
+Performance of PAMSIL clustering Comparison with global upper bound Comparison with constrained upper bound
+Dataset Algorithm $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI $\operatorname{ASW}$ $\texttt{UB}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}(\Delta)}$ $\texttt{UB}_{m}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}_{m}(\Delta)}$
+5x5x5-2 PAMSIL 2 94 $0.883$ $0.834$ $0.669$ $0.899$ $0.256$ $0.715$ $0.064$
+FastMSC 2 94 $0.883$ $0.834$ $0.669$ $0.899$ $0.256$ $0.715$ $0.064$
+DynMSC 8 1 $0.876$ $0.820$ $0.647$ $0.899$ $0.281$ $0.899$ $0.281$
+5x5x5-5 PAMSIL 5 24 $0.675$ $0.751$ $0.478$ $0.899$ $0.468$ $0.762$ $0.373$
+FastMSC 5 19 $0.654$ $0.723$ $0.473$ $0.899$ $0.473$ $0.778$ $0.392$
+DynMSC 2 28 $0.001$ $0.009$ $0.416$ $0.899$ $0.537$ $0.752$ $0.446$
+14x6x6-2 PAMSIL 2 94 $0.883$ $0.834$ $0.640$ $0.879$ $0.271$ $0.695$ $0.078$
+FastMSC 2 92 $0.846$ $0.795$ $0.639$ $0.879$ $0.273$ $0.702$ $0.090$
+DynMSC 9 1 $0.856$ $0.776$ $0.644$ $0.879$ $0.267$ $0.879$ $0.267$
+14x6x6-5 PAMSIL 5 24 $0.670$ $0.739$ $0.473$ $0.888$ $0.467$ $0.752$ $0.371$
+FastMSC 5 23 $0.676$ $0.754$ $0.471$ $0.888$ $0.470$ $0.755$ $0.376$
+DynMSC 6 20 $0.678$ $0.746$ $0.472$ $0.888$ $0.468$ $0.764$ $0.382$
+5.4 Discussion
+In Table 2, we saw that $\texttt{UB}(\Delta)$ was able to confirm that PAMSIL reached a globally optimal $\operatorname{ASW}$ for the synthetic dataset 400-64-2-2.
+This is interesting from a mathematical perspective, but it is not a realistic use case.
+For any real-world dataset, we expect the likelihood of the bound being sharp to be rather low.
+Replacing the standard $\operatorname{ASW}$ range [1,1][-1,1] with the data-dependent interval $1,\texttt{UB}(\Delta)]$ enhances the interpretability of observed $\operatorname{ASW}$ values.
+However, when the empirical $\operatorname{ASW}$ lies far below the upper bound, it remains unclear to what extent further improvement is actually possible.
+We suspect this gap to often be substantial for the global upper bound, even when the empirical $\operatorname{ASW}$ is close to $\operatorname{ASW}(\mathcal{C}^{*},\Delta)$.
+In many real-world applications, however, analysts impose minimum cluster size constraints to avoid overly small or degenerate clusters [24, Höppner2008].
+By incorporating such constraints through $m$, the refined bound $\texttt{UB}_{m}(\Delta)$ better reflects the relevant solution space and provides more meaningful guidance in practice.
+As demonstrated in Table 4, $\texttt{UB}_{m}(\Delta)$ ensures that the PAMSIL solution is within 30 $\%$ of the optimum (in the constrained solution space) for five of the examined datasets.
+The experimental results indicate that the usefulness of the upper bound varies between different contexts; understanding when it is most beneficial in practice is important.
+For the sake of argument, we can assume that the bound is useful whenever it is far from $1.0$ (say, less than $0.5$).
+Following this logic, one can expect the bound to be most informative when the dimensionality of the data is rather high.
+The curse of dimensionality tells us that distances become more similar when dimensionality increases, which should result in the upper bound getting closer to zero.
+Indeed, for the real-world cases, the smallest upper bound we observed was for the Rna dataset with Euclidean distances and 20531 features.
+In contrast, the largest upper bound we observed was for the Customers dataset (Cosine distance), with only 7 features.
+In addition to the fact that the data should be high-dimensional, we speculate that the upper bound is more useful when the number of clusters in the optimal solution is relatively low.
+6 Summary
+6.1 Limitations
+Challenges with the ASW
+A natural limitation of the proposed method stems from the limitations of the silhouette metric itself.
+Many clustering quality indices have been proposed over time, and the ASW is not inherently superior to these.
+Its usefulness depends on the structure of the data.
+For instance, the ASW can be misleading when the underlying clusters are of different diameters or strongly anisotropic, in which case a desirable partition can even result in negative silhouette widths.
+Also, singletons are usually given a silhouette width of 0 by default, even when they are well separated.
+In conclusion, the proposed upper bound should only be considered in situations where the silhouette is meaningful.
+Non-sharpness
+The proposed upper bound is not guaranteed to be close to the maximum ASW.
+Therefore, the upper bound should not be viewed as a precise target, but rather an informative ceiling.
+Generally, we expect the bound to be rather loose when the number of underlying clusters is large.
+Challenges with evaluation
+A fundamental limitation inherent to the problem is that the true global maximum of the ASW is unknown for any non-trivial dataset.
+Consequently, the quality of our upper bound cannot be assessed exactly, but only indirectly through empirical evidence.
+This makes it difficult to draw precise conclusions.
+Poor scalability
+Since the proposed algorithm assumes the full dissimilarity matrix as input, it suffers from the same scalability issue as that for computing the ASW, namely that $\mathcal{O}(n^{2})$ memory is required.
+In principle, one could design an implementation that instead computes and sorts distances per data point, reducing the memory requirement to $\mathcal{O}(n)$.
+However, this does not change the overall time complexity, which remains $\mathcal{O}(n^{2}\log n)$.
+Therefore, the proposed method does not scale to very large datasets.
+Hardware constraints
+Due to hardware constraints, large datasets (e.g., $n>50000$) were not included in our empirical evaluation.
+This does not affect the theoretical correctness of the method, but limits our empirical assessment at large scales.
+6.2 Conclusion
+In this paper, we proposed a dataset-dependent upper bound on the $\operatorname{ASW}$ as a complementary tool for analyzing clustering quality.
+While $\operatorname{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a tighter data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time.
+This upper bound provides a more relevant reference value.
 It often lies well below $1$, reflecting the inherent limitations of the dataset itself.
-Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often nearly tight and yields valuable insights into the structure and limitations of the data.
-At the same time, we emphasize that the bound is not guaranteed to be sharp in general; the true global ASW-maximum may fall considerably below the bound.
-For this reason, the bound should not be viewed as a precise target but as an informative ceiling that frames observed ASW scores in a sharper context.
-From a broader perspective, our main contribution is to establish a proof of concept: that efficiently computable, data-dependent upper bounds on ASW can enrich cluster analysis.
-We believe this idea opens a promising direction for future work, where the focus should be on deriving tighter bounds and extending the framework to other internal clustering validity indices.
-An important next step is to better understand under which conditions the proposed bound is most informative in practice — for instance, how its tightness varies with properties of the data, the dissimilarity metric, the dimensionality of the feature space, or the number of underlying clusters.
+As for the practical usefulness of the proposed upper bound, our findings do not indicate that meaningful insights can be expected in the general case.
+However, it is evident from the results that there are contexts in which the upper bound succeeds in providing useful information.
+For example, for the Rna dataset, we observed an upper bound of $0.42$, which placed the empirical $\operatorname{ASW}$ of $0.23$ in a completely different lighting.
+Pinpointing these cases where the upper bound is the most beneficial would be of great value.
+From a broader perspective, our main contribution has been to establish a proof of concept: that efficiently computable, data-dependent upper bounds on $\operatorname{ASW}$ can enrich cluster analysis.
+We believe that this idea opens a promising direction for future work, where the focus should be on deriving tighter bounds and extending the framework to other internal clustering validity indices, for example, to the medoid-based silhouette.
+An important open theoretical question, which we do not address here, is to characterize how loose the proposed upper bound is in a given case, for example, by deriving a lower bound on the gap between the true maximum $\operatorname{ASW}$ and $\texttt{UB}(\Delta)$.
+As mentioned above, another important next step is to better understand under which conditions the proposed bound is most informative in practice — for instance, how its usefulness varies with properties of the data, the dissimilarity metric, the dimensionality of the feature space, or the number of underlying clusters.
 Such insights would clarify in which contexts the bound provides the greatest practical value.
@@
 CRediT authorship contribution statement
-Hugo Sträng: conceptualization, data curation, methodology, experiment, visualization, writing, editing; Tai Dinh: visualization, writing, editing, validation.
+Hugo Sträng: Conceptualization, Methodology, Investigation, Data Curation, Writing - Original Draft, Writing - Review &\& Editing, Visualization; Tai Dinh: Validation, Writing - Review &\& Editing, Visualization.
 Acknowledgments